### v2

# Improved dimension dependence in the Bernstein von Mises Theorem via a new Laplace approximation bound

Abstract
We prove a new simple and explicit bound on the total variation distance between a measure $\pi\propto e^{-nf}$ on $\mathbb{R}^{d}$ and its Laplace approximation. The bound is proportional to $d/\sqrt{n}$, which has recently been shown to be the tight rate in terms of dimension dependence. Our bound holds under weak regularity conditions on $f$ and at least linear growth of $f$ at infinity. We then apply this bound to prove the first ever Bernstein-von Mises (BvM) theorems on the asymptotic normality of posterior distributions in the regime $n\gg d^{2}$. This improves on the tightest previously known condition, $n\gg d^{3}$. We establish the BvM for the following data-generating models: 1) exponential families, 2) arbitrary probability mass functions on $d+1$ states, and 3) logistic regression with Gaussian design. Our statements of the BvM are nonasymptotic, taking the form of explicit high-probability bounds. We also show in a general setting that the prior can have a stronger regularizing effect than previously known while still vanishing in the large sample limit.
1 Introduction
[…]
$\gamma_{v}:=\mathcal{N}(\hat{\theta},(n\nabla^{2}v(\hat{\theta}))^{-1}),\qquad\hat{\theta}:=\operatorname*{argmin}_{\theta\in\mathbb{R}^{d}}v(\theta).$
[…]
Our first contribution is a bound of this form. To put our bound into context, let us review similar results in the literature. The question of dimension dependence has been studied from different perspectives for several decades, and until recently, $n\gg d^{3}$ has been the best known condition to guarantee an accurate approximation (lowest power of $d$). Earlier works on this topic, which take a more asymptotic and statistical flavor, will be discussed below. More recently, the works [7, 27, 12, 16] have obtained explicit nonasymptotic bounds. The papers [7] and [27] bound the KL divergence, proving an inequality of the form $\mathrm{KL}\left(\,\pi_{v}\;||\;\gamma_{v}\right)\leq c_{3}(v)d^{3}/n$. The works [27] and [12] bound the TV distance, proving an inequality of the form $\mathrm{TV}(\pi_{v},\gamma_{v})\leq c_{3}(v)\sqrt{d^{3}/n}$. (To be more precise, [27]’s bounds are in terms of an effective dimension in place of $d$.) Although the definition of $c_{3}(v)$ varies among these three works, all three definitions involve the tensor norm of the third derivative of $v$ weighted by the Hessian of $v$ at the minimum. Finally, [16] bounds the difference between $\int ge^{-nv}$ and the leading order term in the Laplace asymptotic expansion; note that here, $e^{-nv}$ is an unnormalized density. After suitable rescaling of the error to account for this lack of normalization, the result also scales as $\sqrt{d^{3}/n}$, with a prefactor depending on the second and third derivatives of $v$.
Since then, [13] has proved a very similar bound but, for the first time, with a tighter dimension dependence. After substantial simplification, the authors’ bound can be shown to take the form $\mathrm{TV}(\pi_{v},\gamma_{v})\leq c_{3}(v)\sqrt{d^{2}/n}$; see Section 2.2. The work [14] then showed that $n\gg d^{2}$ is not only sufficient, but in general necessary, by deriving the leading order contribution to the TV error $\mathrm{TV}(\pi_{v},\gamma_{v})$. The leading order term is bounded above by $c_{3}(v)\sqrt{d^{2}/n}$, and an example is provided with a matching lower bound.
We note that most of the bounds described above also contain a negligible term decaying exponentially in $d$ or $n$; we have omitted this term for simplicity.
Although [13, 14] have achieved the optimal dimension dependence, both leave something to be desired. The bound of [13] is quite complicated, and is too coarse as soon as $v$ satisfies mild growth conditions at infinity (see the discussion in Section 2.2 below). The bound of [14] requires that $v\in C^{4}$ and leaves the constant in front of the remainder term unspecified. Our first contribution resolves these issues:
  • We prove an explicit, simple and transparent bound on $\mathrm{TV}(\pi_{v},\gamma_{v})$ assuming $v$ is convex and has a unique global minimizer $\hat{\theta}$.
  • The bound is optimal in its dimension dependence and is tighter than [13] due to the linear growth at infinity of convex functions.
  • We prove our bound under weaker regularity assumptions than those in [13] and [14]: namely, we only require that $v\in C^{2}$, and that $\|\nabla^{2}v(\theta)-\nabla^{2}v(\hat{\theta})\|=\mathcal{O}(\|\theta-\hat{\theta}\|)$ in a small neighborhood of $\hat{\theta}$.
  • We extend the scope of applicability of [13, 14] by allowing for $\pi_{v}$ to be defined on an arbitrary convex domain $\Theta\subseteq\mathbb{R}^{d}$, rather than all of $\mathbb{R}^{d}$.
We make two final notes pertaining to our TV bound. First, convexity of $v$ is not strictly necessary. Rather, only linear growth starting at distance $r\sqrt{d/n}$ away from $\hat{\theta}$ is needed, for some $r>0$. Convexity simply gives the freedom to optimize the bound over $r$. Second, our bound is not specific to Bayesian inference, and can be used in any application where normalized Laplace-type integrals of bounded functions arise.
Bernstein von Mises.
Our second main contribution is specific to Bayesian inference. We therefore now describe in more detail the structure of a posterior measure. The posterior $\pi_{v}$ of the unknown parameter $\theta\in\Theta\subset\mathbb{R}^{d}$ is obtained by using Bayes rule to combine a prior distribution $\pi_{0}$ on $\theta$ with the likelihood $p(Y_{1:n}|\theta)$ of the observed data under the model $Y_{1:n}\sim p(\cdot\mid\theta)$. The result is $\pi_{v}(\theta)\propto p(Y_{1:n}|\theta)\pi_{0}(\theta)$. If $Y_{1:n}=(Y_{1},\dots,Y_{n})$ for independent samples $Y_{i}$, then $p(Y_{1:n}|\theta)$ is a product of $n$ terms, and can therefore be written as $p(Y_{1:n}|\theta)=e^{-n\ell(\theta)}$ for some function $\ell$. Thus $\pi_{v}\propto e^{-n\ell}\pi_{0}=e^{-nv}$ where $v=\ell-\frac{1}{n}\log\pi_{0}$. As $n$ increases, the contribution to $v$ from $n^{-1}\log\pi_{0}$ becomes negligible, and $\ell$ — an average of $n$ independent random variables — converges to its expectation. This justifies our earlier comment that $\pi_{v}\propto e^{-nv}$, where $v$ stabilizes as $n\to\infty$.
Building on our improved Laplace error bound, we tighten the dimension dependence in the Bernstein von Mises Theorem (BvM) for several statistical models. The BvM is a classical result linking Bayesian to frequentist inference: it shows that frequentist confidence regions are asymptotically the same as Bayesian credible sets, so that Bayesian and frequentist inference are two sides of the same coin. In particular, the BvM shows that the prior is negligible in the large sample limit. To be more precise (but informal), the BvM states that when $n\gg 1$, the following approximation holds with high probability under the ground truth data generating distribution $p(\cdot|{\theta^{*}})$ [32]:
$\pi_{v}\approx\gamma^{*}:=\mathcal{N}(\hat{\theta},I({\theta^{*}})^{-1}),\qquad I({\theta^{*}}):=n\mathbb{E}\,[\nabla^{2}\ell({\theta^{*}})].$ (1.1)
Here, ${\theta^{*}}$ is the ground truth parameter and $\hat{\theta}$ is the minimizer of $\ell$ (i.e. maximizer of the likelihood, or MLE). The matrix $I({\theta^{*}})$ is the Fisher information matrix scaled by $n$. Note that here and below, the dependence on $n$ of various quantities (such as $\ell,\hat{\theta}$) is not explicitly indicated.
The BvM is similar to a Laplace approximation error bound. For example, if the prior is flat ($\pi_{0}\equiv 1$) then we can write $\pi_{v}$ as $\pi_{v}=\pi_{\ell}\propto e^{-n\ell}$. In this case, the Laplace approximation to $\pi_{v}=\pi_{\ell}$ is $\pi_{\ell}\approx\gamma_{\ell}=\mathcal{N}(\hat{\theta},(n\nabla^{2}\ell(\hat{\theta}))^{-1})$. The Gaussian $\gamma^{*}$ from (1.1) is very similar to $\gamma_{\ell}$: the means of the two Gaussians are the same, and the inverse covariances $I({\theta^{*}})=n\mathbb{E}\,[\nabla^{2}\ell({\theta^{*}})]$ and $n\nabla^{2}\ell(\hat{\theta})$ are close for $n$ large. Beyond this slight difference in definition, the more significant difference is that the BvM considers the posterior as a random object which varies with random draws of the data. In the BvM, the statement that $\pi_{v}\approx\gamma^{*}$ is shown to hold with high probability over these random draws. Thus the BvM comes with a stronger guarantee: the posterior is approximately Gaussian for large $n$, regardless of which data is drawn. In contrast, upper bounds on the Laplace approximation error are typically expressed in terms of a fixed dataset.
The works [9, 10] by Ghosal were the first to prove the BvM in the regime in which dimension $d$ grows large with $n$. The settings Ghosal considered were posteriors arising from linear regression models and exponential families, respectively. The exponential family BvM was later tightened and generalized in [5]. The work [6] proves a BvM for the posterior of a discrete probability mass function truncated to its first $d$ entries. [26] proves a BvM for parametric models under more general conditions, which include e.g. generalized linear models (GLMs). [18] proves a BvM for nonlinear Bayesian inverse problems. Up to model-specific factors, each of these works requires at least that $n\gg d^{3}$. See these works for further references on BvMs with $d$ growing. There is also a vast literature on the BvM in fixed dimension; see [21] and the references therein.
In high dimensions $d$, the condition $n\gg d^{3}$ on sample size is very restrictive, and seemed to put Bayesian inference at a disadvantage compared to frequentist inference. To enable tractable computation, both statistical frameworks rely on asymptotic normality: of the posterior in Bayesian inference, and of the MLE in frequentist inference. Yet in frequentist inference, the pioneering work [22] showed that for exponential families, the MLE is asymptotically normal provided $d^{2}/n\to 0$[11] extended this result to more general parametric models, while [25] proved finite sample error bounds on the Fisher expansion of the MLE (closely related to normality), also provided $n\gg d^{2}$.
In this work we put Bayesian and frequentist inference on equal footing:
  • We prove for the first time that the BvM holds when $n\gg d^{2}$ for the following common statistical models (listed are the data-generating distribution $p(\cdot|\theta)$, $\theta\in\Theta\subset\mathbb{R}^{d}$):
    1. 1.
      exponential families in canonical form,
    2. 2.
      an arbitrary probability mass function on $d+1$ states,
    3. 3.
      logistic regression with Gaussian design.
  • In each case, we show that $\mathrm{TV}(\pi_{v},\gamma^{*})\leq c_{3}\sqrt{d^{2}/n}$ with high probability (up to negligible terms) where $c_{3}$ is now a deterministic, model dependent coefficient.
  • We show that the prior $\pi_{0}$ is negligible when $\|\nabla\log\pi_{0}\|=o(\sqrt{n})$, relaxing previous conditions on the prior from [9, 10, 5, 18, 26] by a factor of $\sqrt{d}$.
  • For the exponential family model, we show that if the family is log concave, then $c_{3}$ is an absolute constant.
The dimension dependence in these three BvMs improves on [10, 5] for the first model, [6] for the second model, and [26] for the third model. Our relaxed condition on the prior is also notable: it quantifies more precisely how strong of a regularizing effect the prior can have while still being negligible in the large sample limit. See Section 3.3 for a more precise statement of our conditions on the prior compared to previous works.
We take note of two more features of our BvM results. First, our work improves on most previous results by making the statement of the BvM nonasymptotic. We give explicit expressions (up to absolute constants) for both the upper bound and the probability with which it holds, in terms of fixed $d$ and $n$. To our knowledge, only [26] has formulated the BvM nonasymptotically.
Secondly, in our BvM for the logistic regression model, the high probability bound is with respect to the randomness of both the label distribution and the Gaussian feature distribution. Our treatment of the random design case seems to be new; we are not aware of other BvM results for linear regression or generalized linear models (GLMs) with random design.
The key tool we use in the BvM proofs is our Laplace approximation error bound. To deduce the BvM from this bound, it essentially remains to prove concentration result for various i.i.d. averages (or suprema thereof). Most challenging of the required concentration results is a high probability bound on the operator norm of the third derivative tensor of the log likelihood. These additional probabilistic steps in the BvM proof are model dependent and have been addressed in prior works on the BvM, some of which have been cited above. We hope that our Laplace approximation bound can be combined with these prior results to prove the BvM with tighter dimension dependence in a wider range of models.
Other related works.
[…]
We also mention a few works which prove high probability bounds on the Laplace approximation error $\mathrm{err}(\pi_{v},\gamma_{v})$. This is contrast to the Laplace bounds cite above, which are fixed-sample. In [3], the authors prove that for sparse generalized linear models with $n$ samples and $q$ active covariates, the relative Laplace approximation error of the posterior normalizing constant is bounded by $(q^{3}\log^{3}n/n)^{1/2}$ with high probability. In [30], the authors study the pointwise ratio between the density of the posterior $\pi_{v}$ and that of the Laplace approximation $\gamma_{v}$. The authors show that for generalized linear models with Gaussian design, the relative error is bounded by $d^{3}\log n/n$ with probability tending to 1, provided $d^{3+\epsilon}\ll n$ for some $\epsilon>0$. For logistic regression with Gaussian design, they obtain the improved rate $d^{2}\log n/n$ provided $d^{2.5}\ll n$.
Finally, in addition to obtaining a fixed-sample Laplace bound in terms of $d^{3}/n$, the aforementioned work [7] also derives a high probability bound. The author shows $\mathrm{KL}\left(\,\pi_{v}\;||\;\gamma_{v}\right)$ goes to 0 with probability tending to 1 if $d^{3}/n\to 0$.
Organization
The rest of the paper is organized as follows. In Section 2 we state, discuss, and outline the proof of our new Laplace approximation bound in a general setting, where $\pi_{v}\propto e^{-nv}$ need not have the structure of a posterior. In Section 3 we introduce the Bayesian setting and use our Laplace bound to prove some key preliminary results for the BvM proofs. In Sections 456 we prove the BvM for posteriors stemming from exponential families, observations of a discrete probability distribution, and a logistic regression model, respectively. Omitted proofs can be found in the Appendix.
Notation
The letter $C$ always denotes an absolute constant. The inequality $a\lesssim b$ denotes $a\leq Cb$ for an absolute constant $C$. We let ||$\cdot$|| ⋅ | denote the standard Euclidean norm in $\mathbb{R}^{d}$. If $A\in\mathbb{R}^{d\times d}$ is a symmetric positive definite matrix and $u\in\mathbb{R}^{d}$ is a vector, we let
$|u|_{A}=|A^{1/2}u|.$
Let $k\geq 1$ and $S$ be a symmetric $k$-linear form, identified with its symmetric tensor representation. We define the $A$-weighted operator norm of $S$ to be
$\|S\|_{A}:=\sup_{|u|_{A}=1}\langle S,u^{\otimes k}\rangle=\sup_{|u|_{A}=1}\sum_{i_{1},\dots,i_{k}=1}^{d}S_{i_{1}i_{2}\dots i_{k}}u_{i_{1}}u_{i_{2}}\dots u_{i_{k}}$ (1.2)
[…]
$\|S\|_{A}=|A^{-1/2}S|,\qquad|S|_{A}=|A^{1/2}S|.$
Thus it is important to distinguish between the meaning of $\|\cdot\|_{A}$ and $|\cdot|_{A}$ when the quantity inside the norm is a vector. By Theorem 2.1 of [34], for symmetric tensors the definition (1.2) coincides with the standard definition of operator norm:
$\sup_{|u_{1}|_{A}=\dots=|u_{k}|_{A}=1}\langle S,u_{1}\otimes\dots\otimes u_{k}\rangle=\|S\|_{A}=\sup_{|u|_{A}=1}\langle S,u^{\otimes k}\rangle.$
2 Laplace approximation bound
We begin by stating and discussing our general bound on the Laplace approximation accuracy. In Section 2.1, we outline the proof. In Section 2.2, we compare our bound to other bounds in the literature which achieve the tight $d/\sqrt{n}$ rate. We consider a general setting which need not stem from Bayesian inference. Throughout the paper, we let $\Theta$ denote a convex subset of $\mathbb{R}^{d}$. The value $n$ should be thought of as some large positive number.
Definition 2.1.
Let $f\in C^{2}(\Theta)$. Whenever $e^{-nf}$ is integrable, we define the probability density $\pi_{f}(\theta)\propto e^{-nf(\theta)}$, $\theta\in\Theta$. If $f$ has a unique strict global minimizer $\hat{\theta}\in\Theta$, we define the following Laplace approximation to $\pi_{f}$:
$\gamma_{f}=\mathcal{N}(\hat{\theta},(nH)^{-1}),\qquad H:=\nabla^{2}f(\hat{\theta}).$
Theorem 2.2.
Let $f\in C^{2}(\Theta)$ be convex with unique strict global minimizer $\hat{\theta}$. Assume there is $r\geq 6$ such that ${\mathcal{U}}(r)\subset\Theta$, where ${\mathcal{U}}(r)=\{|\theta-\hat{\theta}|_{H}\leq r\sqrt{d/n}\}$. Let
$\delta_{3}(r)=\sup_{\theta\in{\mathcal{U}}(r)}\frac{\|\nabla^{2}f(\theta)-\nabla^{2}f(\hat{\theta})\|_{H}}{|\theta-\hat{\theta}|_{H}}$ (2.1)
and suppose $r\delta_{3}(r)\sqrt{d/n}\leq 1/2$. Then $e^{-nf}\in L^{1}(\Theta)$ so $\pi_{f}\propto e^{-nf}$ is well-defined, and we have
$\mathrm{TV}(\pi_{f},\;\gamma_{f})\leq\frac{\delta_{3}(r)d}{\sqrt{2n}}+3e^{-dr^{2}/9}.$ (2.2)
Remark 2.3 (Convexity assumption).
To prove (2.2), we use convexity only through the linear growth lower bound (2.6) in Lemma 2.8 below. Therefore, convexity of $f$ can be replaced by (2.6) in the statement of Theorem 2.2. However, under convexity (2.6) holds for any $r\geq 0$, giving us the flexibility to optimize the bound (2.2) over $r$.
Remark 2.4 (Smoothness and $\delta_{3}$).
As noted in the introduction, the only smoothness we require is that $f\in C^{2}(\Theta)$ and $\|\nabla^{2}f(\theta)-\nabla^{2}f(\hat{\theta})\|_{H}=\mathcal{O}(|\theta-\hat{\theta}|_{H})$ for $\theta\in{\mathcal{U}}(r)$, to ensure that $\delta_{3}(r)<\infty$. Thus for example $f(\theta)=|\theta|^{2}/2+|\theta|^{3}/6$ satisfies the conditions of the theorem: it is strictly minimized at $\hat{\theta}=0$ and $f\in C^{2}$, with
$\nabla^{2}f(\theta)=\left(1+\frac{1}{2}|\theta|\right)I_{d}+\frac{1}{2|\theta|}\theta\theta^{\intercal},\quad\theta\neq 0$
and $\nabla^{2}f(0)=I_{d}$. We therefore have $\delta_{3}(r)\leq 1$ for all $r>0$, even though $\nabla^{3}f$ does not exist at $\theta=0$. If $f$ is $C^{3}$ in a neighborhood of $\hat{\theta}$, then one can bound $\delta_{3}(r)$ via
$\delta_{3}(r)\leq\sup_{\theta\in{\mathcal{U}}(r)}\|\nabla^{3}f(\theta)\|_{H}$ (2.3)
However, even when the third derivative tensor exists, the original definition (2.1) of $\delta_{3}$ may in some cases be easier to compute, because it avoids the use of tensor operator norms.
Remark 2.5 (Choice of $r$).
Note that (2.2) is valid for all $r\geq 6$ such that ${\mathcal{U}}(r)\subseteq\Theta$ and $r\delta_{3}(r)\sqrt{d/n}\leq 1/2$. Thus $r$ can be chosen to optimize the bound. If $d$ is very large, then the second term in the bound (2.2) is exponentially small in $d$, so it suffices to choose $r=6$. If $d$ cannot be considered too large, then the optimal choice of $r$ depends on how $\delta_{3}(r)$ scales with $r$ and $d$. The following example shows how to choose $r$ in the special case when $\delta_{3}(r)$ is uniformly bounded.
Example 2.6 (Uniformly bounded third derivative).
If $f\in C^{3}(\Theta)$ and $\sup_{\theta\in\Theta}\|\nabla^{3}f(\theta)\|_{H}\leq C_{3}$ then
$\mathrm{TV}(\pi_{f},\;\gamma_{f})\leq C_{3}\left(\frac{d}{\sqrt{2n}}+\frac{9}{\sqrt{n}}\right)$ (2.4)
for all $d/n$ small enough. Indeed, note that $C_{3}$ is a uniform bound on $\delta_{3}(r)$ over all $r\geq 0$, so (2.2) is optimized by taking $r$ as large as possible while satisfying the constraint. In other words, take $r=\sqrt{n/d}/2C_{3}$. This choice of $r$ satisfies $r\geq 6$ provided $\sqrt{n/d}\geq 12C_{3}$. Substituting $r=\sqrt{n/d}/2C_{3}$ and $\delta_{3}(r)\leq C_{3}$ into (2.2) gives $\mathrm{TV}(\pi_{f},\;\gamma_{f})\leq\frac{1}{\sqrt{2}}C_{3}d/\sqrt{n}+3\,\mathrm{exp}(-n/(36C_{3}^{2}))$. Applying the inequality $e^{-x^{2}/4}\leq 1/x$, $x>0$ with $x=\sqrt{n}/(3C_{3})$ yields (2.4).
Remark 2.7 (Affine invariance).
If $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is a bijective affine map, then the pushforward $T_{\#}\pi_{f}$ of $\pi_{f}$ under $T$ is given by $T_{\#}\pi_{f}=\pi_{f\circ T^{-1}}$. Also, one can check that $\gamma_{f\circ T^{-1}}=T_{\#}\gamma_{f}$. Since the TV distance is invariant under bijective coordinate transformations, we therefore have $\mathrm{TV}(\pi_{f},\gamma_{f})=\mathrm{TV}(T_{\#}\pi_{f},T_{\#}\gamma_{f})=\mathrm{TV}(\pi_{f\circ T^{-1}},\gamma_{f\circ T^{-1}}).$ Due to this invariance, a good upper bound on $\mathrm{TV}(\pi_{f},\gamma_{f})$ should also be affine invariant, and so should the conditions under which the bound holds. It is straightforward to check that Theorem 2.2 has this property.
2.1 Proof outline
The assumptions of Theorem 2.2 imply two key properties used in the proof: first, $\pi_{f}$ is strongly log concave in the neighborhood ${\mathcal{U}}(r)$ of $\hat{\theta}$. Second, $f$ grows linearly away from $\hat{\theta}$ in ${\mathcal{U}}(r)^{c}$. These two observations are stated in the following lemma:
Lemma 2.8 (Hessian lower bound and linear growth).
Under the assumptions of Theorem 2.2, it holds
$\nabla^{2}f(\theta)\succeq\frac{1}{2}H,\quad\forall\theta\in{\mathcal{U}}(r).$ (2.5)
Furthermore, we have the following growth bound:
$f(\theta)-f(\hat{\theta})\geq\frac{r}{4}\sqrt{d/n}|\theta-\hat{\theta}|_{H}\qquad\forall\theta\in\Theta\setminus{\mathcal{U}}(r).$ (2.6)
See Appendix A for the proof of the lemma. The bound (2.5) does not require convexity. Similarly, (2.6) only uses convexity via the property of convex functions that the infimum of $(f(\theta)-f(\hat{\theta}))/|\theta-\hat{\theta}|_{H}$ over $\theta\in{\mathcal{U}}(r)^{c}$ is achieved on the boundary $\{|\theta-\hat{\theta}|_{H}=r\sqrt{d/n}\}$. Thus any other function with this property will satisfy the lemma. Note also that (2.6) proves $e^{-nf}\in L_{1}(\Theta)$.
Next, the following lemma presents the main tools in the proof of Theorem 2.2.
Lemma 2.9.
Let $\mu\propto F$ and $\hat{\mu}\propto\hat{F}$ be two probability densities on $\mathbb{R}^{d}$ and let ${\mathcal{U}}\subset\mathbb{R}^{d}$ be such that $\mu({\mathcal{U}})>0,\hat{\mu}({\mathcal{U}})>0$. Let $\mu|_{\mathcal{U}}$, $\hat{\mu}|_{\mathcal{U}}$ be the restrictions of $\mu$ and $\hat{\mu}$ to ${\mathcal{U}}$, respectively. Then
$\mathrm{TV}(\mu,\hat{\mu})\leq\frac{\int_{{\mathcal{U}}^{c}}F}{\int F}+\frac{\int_{{\mathcal{U}}^{c}}\hat{F}}{\int\hat{F}}+\mathrm{TV}\left(\mu|_{\mathcal{U}},\hat{\mu}|_{\mathcal{U}}\right)$ (2.7)
Furthermore, suppose $\mu$ and $\hat{\mu}$ are strictly positive on ${\mathcal{U}}$ with $\mu\propto e^{-nf}$ and $\hat{\mu}\propto e^{-n\hat{f}}$ where $f\in C^{2}({\mathcal{U}})$, $\hat{f}\in C^{1}({\mathcal{U}})$. Suppose also that $\nabla^{2}f(\theta)\succeq\lambda H$ for all $\theta\in{\mathcal{U}}$ and some $H\succ 0$. Then
$\mathrm{TV}\left(\mu|_{\mathcal{U}},\hat{\mu}|_{\mathcal{U}}\right)^{2}\leq\frac{n}{4\lambda}\mathbb{E}\,_{X\sim\hat{\mu}|_{\mathcal{U}}}\left[\|\nabla(f-\hat{f})(X)\|_{H}^{2}\right],$ (2.8)
and therefore the total TV distance between $\mu$ and $\hat{\mu}$ is bounded as
$\mathrm{TV}(\mu,\hat{\mu})\leq\frac{\int_{{\mathcal{U}}^{c}}e^{-nf}}{\int e^{-nf}}+\frac{\int_{{\mathcal{U}}^{c}}e^{-n\hat{f}}}{\int e^{-n\hat{f}}}+\sqrt{\frac{n}{4\lambda}}\mathbb{E}\,_{X\sim\hat{\mu}|_{\mathcal{U}}}\left[\|\nabla(f-\hat{f})(X)\|_{H}^{2}\right]^{\frac{1}{2}}.$ (2.9)
See Appendix A.1 for the proof of this lemma. The bound (2.8) stems from a log Sobolev inequality (LSI) [2], which can be applied since $\mu$ is strongly log-concave in ${\mathcal{U}}$. The inspiration for this technique comes from [13], who also prove a bound on $\mathrm{TV}(\pi_{f},\gamma_{f})$ in the context of Bayesian inference. Using an LSI offers a very simple route to the tighter dimension dependence, as discussed in [14, Section 4.2]. See Section 2.2 for a comparison between our bound and that of [13].
We apply (2.9) with ${\mathcal{U}}={\mathcal{U}}(r)$, $\mu=\pi_{f}$ (extended to be zero in $\mathbb{R}^{d}\setminus\Theta$) and $\hat{\mu}=\gamma_{f}\propto e^{-n\hat{f}}$, where $\hat{f}(\theta)=\frac{1}{2}|\theta-\hat{\theta}|^{2}_{H}$. Since $\nabla^{2}f\succeq\frac{1}{2}H$ on ${\mathcal{U}}(r)$ by (2.5), the lower bound condition in the lemma is satisfied with $\lambda=1/2$. Therefore, in our setting, the bound (2.9) takes the form
$\mathrm{TV}(\pi_{f},\gamma_{f})\leq\frac{\int_{\Theta\setminus{\mathcal{U}}}e^{-nf}}{\int_{{\mathcal{U}}}e^{-nf}}+\gamma_{f}({\mathcal{U}}^{c})+\sqrt{\frac{n}{2}}\mathbb{E}\,_{\theta\sim\gamma_{f}|{\mathcal{U}}}\left[\|\nabla f(\theta)-H(\theta-\hat{\theta})\|^{2}_{H}\right]^{\frac{1}{2}}.$ (2.10)
We see that to bound the TV distance between $\pi_{f}$ and $\gamma_{f}$, it suffices to bound the two tail integrals and the local expectation in (2.10). The first tail integral can be bounded using the linear growth of $f$. The bound on $\gamma_{f}({\mathcal{U}}^{c})$ follows by standard Gaussian tail inequalities. Finally, to bound the local expectation we use a Taylor expansion of $\nabla f(\theta)$ about $\hat{\theta}$ to get that $\nabla f(\theta)-H(\theta-\hat{\theta})=(\nabla^{2}f(\xi)-\nabla^{2}f(\hat{\theta}))(\theta-\hat{\theta})$ for a point $\xi$ between $\theta$ and $\hat{\theta}$. From here we use the definition of $\delta_{3}$ and simple Gaussian calculations show that the third term in (2.10) is bounded as $\delta_{3}(r)d/\sqrt{n}$. See Appendix A.2 for these calculations, which finish the proof of Proposition 2.2.
2.2 Comparison with [14] and [13].
A nice property of the bound (2.2) in Proposition 2.2 is its remarkable simplicity, which we hope will lead to broad applicability. So far, only two other works have obtained bounds on $\mathrm{TV}(\pi_{f},\gamma_{f})$ scaling as $d/\sqrt{n}$ — that of [14] and [13].
The work [14] decomposes the TV distance into a leading order term and a negligible remainder. In the $d/\sqrt{n}\ll 1$ regime, the leading order term gives precise control on the error. However, there are two important downsides: first, the result requires $f\in C^{4}$. Second, the remainder term in the error bound is only computed up to an absolute constant. Formally, therefore, the result does not give an explicit upper bound.
The work [13] considers the Bayesian setting, where $\pi_{f}$ can be decomposed as $\pi_{f}\propto e^{-n\ell(\theta)}\pi_{0}(\theta)$. Here, $\pi_{0}$ is a proper prior, meaning, $\pi_{0}$ is a probability density integrating to 1. As a result of explicitly incorporating this structure, the bound in [13] is significantly more complex than ours. It involves multiple quantities: $\bar{M}_{2},\bar{J}_{n}(\bar{\theta}_{n}),\bar{\delta}$ (analogous to our $\delta_{3}(r),H,r$), but also five additional quantities: $M_{2},\hat{J}_{n}(\hat{\theta}_{n}),\delta,\hat{M}_{1},\kappa$ defined in terms of the likelihood and prior.
Nevertheless, there is one case in which our two bounds can be explicitly compared: when the prior is uniform over a bounded convex subset $\Theta$ of $\mathbb{R}^{d}$, i.e. $\pi_{0}=\frac{1}{\mathrm{Vol}(\Theta)}\mathds{1}_{\Theta}$. Thus $\pi_{f}\propto e^{-nf}$ on $\Theta$, where $f=\ell$. In this setting, we now state the bound of [13], changing their notation slightly:
$\begin{split}\mathrm{TV}(\pi_{f},\gamma_{f})\leq\frac{\mathrm{Tr}(H_{f}^{-1})M_{f}}{4\sqrt{(\lambda_{\min}(H_{f})-rM_{f}\sqrt{d/n})(1-\mathscr{D}_{f})}\sqrt{n}}+2\mathscr{D}_{f}+2\frac{\sqrt{\det\left(H_{f}^{p}\right)}\mathrm{Vol}(\Theta)}{(2\pi)^{d/2}(1-\mathscr{D}_{f}^{p})}n^{d/2}e^{-n\kappa_{f}}\end{split}$ (2.11)
Here,
$\begin{split}H_{f}&=\nabla^{2}f(\hat{\theta}),\\ H_{f}^{p}&=H_{f}+\frac{M_{f}}{3}r\sqrt{d/n}I_{d},\\ M_{f}&=\sup_{|\theta-\hat{\theta}|\leq r\sqrt{d/n}}\|\nabla^{3}f(\theta)\|,\\ \kappa_{f}&=\inf_{|\theta-\hat{\theta}|\geq r\sqrt{d/n}}f(\theta)-f(\hat{\theta}),\\ \mathscr{D}_{f}&=\,\mathrm{exp}\left(-\frac{1}{2}\left(r\sqrt{d}-\sqrt{\mathrm{Tr}(H_{f}^{-1})}\right)^{2}\lambda_{\min}(H_{f})\right),\\ \mathscr{D}_{f}^{p}&=\,\mathrm{exp}\left(-\frac{1}{2}\left(r\sqrt{d}-\sqrt{\mathrm{Tr}((H_{f}^{p})^{-1})}\right)^{2}\lambda_{\min}(H_{f}^{p})\right),\end{split}$ (2.12)
and $\hat{\theta}$ is the unique global minimizer of $f$ on $\Theta$. The bound (2.11) is not affine invariant; for example, if $f$ takes the form $f(\theta)=\epsilon^{2}\theta_{1}^{2}/2+g(\theta_{2},\dots,\theta_{d})$, $\epsilon\ll 1$, then $\lambda_{\min}(H_{f})=\epsilon$ will appear in the denominator of the first term in (2.11), causing the bound to blow up. In contrast, our bound stays fixed as $\epsilon\to 0$, since $\epsilon$ can be removed by a linear rescaling of space. However, this is not entirely fair. Indeed, precisely because of the affine invariance of TV distance (recall Remark 2.7) we can first change variables $f\mapsto f\circ T^{-1}$ in the most favorable way possible. Specifically, we choose a rescaling which transforms the Hessian $H_{f}$ into the identity matrix. This gives rise to the following bound; see Appendix A.3 for details.
$\mathrm{TV}(\pi_{f},\gamma_{f})\leq C\delta_{3}(r)\frac{d}{\sqrt{n}}+2\,\mathrm{exp}\left(-\frac{d}{2}(r-1)^{2}\right)+2C^{d}\mathrm{Vol}(H_{f}^{1/2}\Theta)n^{d/2}e^{-n\kappa},$ (2.13)
where
$\kappa=\inf_{|\theta-\hat{\theta}|_{H_{f}}\geq r\sqrt{d/n}}f(\theta)-f(\hat{\theta}),$ (2.14)
and we have used the upper bound in (2.3) as the definition of $\delta_{3}$ in (2.13). One can check that this bound is now affine invariant. We make one more simplification before discussing the bound. To get (2.13) from (2.11), we not only rescaled $f$ but also made the assumption that $r\delta_{3}(r)\sqrt{d/n}<1$, which is necessary in order for the bound not to blow up. This assumption can be used to show $f(\theta)-f(\hat{\theta})$ is at most $r^{2}d/n$ for all $|\theta-\hat{\theta}|_{H_{f}}=r\sqrt{d/n}$; see Appendix A.3 for this argument. Therefore, $\kappa\leq r^{2}d/n$. This largest possible $\kappa$ leads to the smallest possible righthand side of (2.13). Assuming the best case scenario that $\kappa=r^{2}d/n$, the bound (2.13) further simplifies to
$\mathrm{TV}(\pi_{f},\gamma_{f})\leq C\delta_{3}(r)\frac{d}{\sqrt{n}}+2\,\mathrm{exp}\left(-\frac{d}{2}(r-1)^{2}\right)+2C^{d}\mathrm{Vol}(H^{1/2}\Theta)n^{d/2}e^{-r^{2}d}.$ (2.15)
The first term in (2.15) is the same as our first term in (2.2). This is to be expected: as we have mentioned, our approach to bounding the error in the neighborhood ${\mathcal{U}}(r)=\{\theta\;:|\theta-\hat{\theta}|_{H}\leq r\sqrt{d/n}\}$ of $\hat{\theta}$ (via an LSI) is inspired by that of [13]. The only difference is that we apply the LSI in an affine invariant way. Similarly, the second term of (2.15) is essentially the same as our second term in (2.2).
But the third term in (2.15) is not present in our bound, because we make stronger assumptions on $f$. Namely, we assume $f$ grows at least linearly away from $\hat{\theta}$ in ${\mathcal{U}}^{c}$. In contrast, [13] assumes only that $f-f(\hat{\theta})>\kappa>0$ in ${\mathcal{U}}(r)^{c}$. Thus it could be that $f$ is quadratic in ${\mathcal{U}}(r)$ and flat outside of ${\mathcal{U}}(r)$. If the domain $\Theta$ is sufficiently large, then the flat region of $\pi_{f}$ will eventually dominate: the third term of (2.15) will become large, and $\pi_{f}$ will be far from Gaussian.
Of course, one can compensate for a large $\mathrm{Vol}(H^{1/2}\Theta)$ by choosing $r$ to be large, but this makes the first, leading order term $\delta_{3}(r)d/\sqrt{n}$ large. In fact, whether or not $\mathrm{Vol}(H^{1/2}\Theta)$ is large, $r$ must be at least large enough to ensure $n^{d/2}e^{-r^{2}d}$ is small. Indeed, suppose for simplicity that $\mathrm{Vol}(H^{1/2}\Theta)=1$. In order for the third term of (2.15) to be small, we must have that $r^{2}\gtrsim\log n$. In contrast, in order for the second term $e^{-dr^{2}/9}$ in (2.2) to be small relative to the first term, it suffices that $r^{2}\gtrsim\frac{\log n}{d}\vee 1$ (the maximum with 1 is needed to satisfy our requirement that $r\geq 6$).
For concreteness, let us consider the case $\log n\lesssim d$, so that it suffices to take $r\sim 1$ in our bound (2.2). Let us now consider the effect of $r$ on the leading order term $\delta_{3}(r)d/\sqrt{n}$ of both bounds. We use the following approximate bound on $\delta_{3}(r)$:
$\delta_{3}(r)\lesssim\|\nabla^{3}f(\hat{\theta})\|_{H}+r\sqrt{d/n}\|\nabla^{4}f(\hat{\theta})\|_{H}=\|\nabla^{3}f(\hat{\theta})\|_{H}\left(1+r\sqrt{d/n}\frac{\|\nabla^{4}f(\hat{\theta})\|_{H}}{\|\nabla^{3}f(\hat{\theta})\|_{H}}\right)$ (2.16)
This shows that $r$ starts to play a role in the bounds (2.15) and (2.2) as soon as
$\alpha:=r\sqrt{d/n}\|\nabla^{4}f(\hat{\theta})\|_{H}/\|\nabla^{3}f(\hat{\theta})\|_{H}\gtrsim 1.$
Now, suppose for example that $d/\sqrt{n}$ is a small constant, and that $\|\nabla^{4}f(\hat{\theta})\|_{H}/\|\nabla^{3}f(\hat{\theta})\|_{H}\sim\sqrt{d}$ (the fourth derivative tensor norm growing faster with $d$ than the third derivative tensor norm is not uncommon). Then $\alpha\sim d/\sqrt{n}$ when $r\sim 1$, but $\alpha\sim\sqrt{\log n}(d/\sqrt{n})$ when $r\sim\sqrt{\log n}$. Thus a larger $r$ could push $\alpha$ over the edge from bounded to unbounded.
In summary, when the linear growth assumption is satisfied, it is preferable to use (2.2) over (2.15).
3 Applications to Bayesian inference
Theorem 2.2 naturally applies in the Bayesian context, where approximation of posterior distributions by Gaussians is both of practical and theoretical interest. In Section 3.1 we review the general Bayesian set-up and discuss how Gaussian approximation of posteriors is used given observed (fixed) data, and how Theorem 2.2 can be used to quantify the accuracy of the approximation. In Section 3.2 we introduce the BvM, the “high-probability” counterpart to the fixed sample bound. Proving the BvM holds with rate $d^{2}/n$ is our main application of Theorem 2.2. In Section 3.3 we use Theorem 2.2 to prove the key preliminary results, which we use in Sections 456 to finish the BvM proof for our three statistical models.
3.1 Bayesian preliminaries and Laplace approximation to posteriors
Recall from the introduction that the posterior $\pi_{v}$ takes the form $\pi_{v}(\theta)\propto p(Y_{1:n}\mid\theta)\pi_{0}(\theta)$, where $Y_{1:n}$ is the observed data, $p(\cdot|\theta)$, $\theta\in\Theta\subset\mathbb{R}^{d}$ is the parameterized family of data distributions, and $\pi_{0}$ is the prior. We assume the support of $p$ does not vary with $\theta$. This implies that if $Y_{1:n}$ is indeed drawn from $p(\cdot\mid{\theta^{*}})$ for a ground truth ${\theta^{*}}\in\Theta$, then the likelihood $L(\theta):=p(Y_{1:n}\mid\theta)$ is strictly positive for all $\theta\in\Theta$. We can therefore write $L(\theta)=e^{-n\ell(\theta)}$, where $\ell:\Theta\to\mathbb{R}$ can depend on $n$. We summarize these concepts and introduce a few more in the following definition.
Definition 3.1.
Without loss of generality, redefine $\Theta$ to be its intersection with the support of the prior $\pi_{0}$, so that
$\pi_{0}(\theta)\propto e^{-v_{0}(\theta)},\quad\theta\in\Theta.$
The likelihood is the function
$L(\theta)=p(Y_{1:n}\mid\theta)\propto e^{-n\ell(\theta)},\quad\theta\in\Theta$
where $\ell$ is the negative normalized log likelihood $\ell=-\frac{1}{n}\log L$. We assume $Y_{1:n}\sim p(\cdot\mid{\theta^{*}})$ for a ground truth ${\theta^{*}}\in\Theta$, and we define the negative population log likelihood to be the function
$\ell_{\infty}(\theta)=\mathbb{E}\,_{Y_{1:n}\sim p(\cdot\mid{\theta^{*}})}[\ell(\theta)].$
The posterior distribution is the probability density $\pi_{v}$ such that
$\pi_{v}(\theta)\propto e^{-n\ell(\theta)}\pi_{0}(\theta)\propto e^{-nv(\theta)},\qquad v=\ell+n^{-1}v_{0},\quad\theta\in\Theta.$
Making the Laplace approximation $\gamma_{v}$ to $\pi_{v}$ allows us to compute integrals against the posterior using
[…]
As noted in the introduction, the Gaussian integral is much easier to compute: if $g$ is a polynomial then the integral can be evaluated exactly; otherwise, Gaussian sampling can be used. Importantly, obtaining the Laplace approximation $\gamma_{v}$ only requires us to find the minimizer of $v$ and to compute the Hessian of $v$ at this point. Neither the normalizing constant of $\pi_{v}$ nor any other integrals against $e^{-nv}$ need to be computed. Moreover, the function $v$ is available to the practitioner: it is entirely determined from the statistical model $p(\cdot\mid\theta)$, the prior $\pi_{0}$, and the observed data $Y_{1:n}$.
In the fixed-sample setting, one studies the accuracy of the approximation (3.1) for a given fixed dataset $Y_{1:n}$. As long as $v=\ell+n^{-1}v_{0}$ is convex and $C^{3}$, and has a unique global minimum, Theorem 2.2 can be applied to bound the approximation error. Up to the exponentially negligible term in (2.2), we obtain the bound $\mathrm{TV}(\pi_{v},\gamma_{v})\lesssim\delta_{3}(r)d/\sqrt{n}$. Note that the righthand side depends on the data $Y_{1:n}$ via $\delta_{3}(r)$, since this coefficient is expressed in terms of derivatives of $v$, and $v$ depends on the data.
Applying (2.2) to bound the Laplace approximation accuracy to a posterior seems deceptively simple, but there are challenges in applying the formula, particularly computing $\delta_{3}$. Since fixed sample Laplace accuracy bounds are not our focus here, we refer to the reader to [13] for a discussion of practical and numerical considerations in e.g. choosing a radius $r$ and computing the operator norm of the third order derivative tensor.
3.2 BvM: statement and discussion
We now move on to the BvM, our main application of Theorem 2.2. Classically, the BvM states that $\mathrm{TV}(\pi_{v},\gamma^{*})=o(1)$ with probability tending to 1 as $n\to\infty$ [32]. Here,
$\gamma^{*}=\mathcal{N}(\hat{\theta},(n\nabla^{2}\ell_{\infty}({\theta^{*}}))^{-1}),\qquad\text{where}\quad\hat{\theta}=\operatorname*{argmin}_{\theta\in\Theta}\ell(\theta).$
Note that $\hat{\theta}$ is the MLE (maximizer of the likelihood). The probability is with respect to $Y_{1:n}\sim p(\cdot|{\theta^{*}})$, the ground truth data distribution. For the significance and interpretation of the BvM, see the introduction.
Here, we will be interested in a nonasymptotic version of the BvM. Namely, we will prove a result of the form: “under certain conditions on $d$, $n$, and the statistical model, $\mathrm{TV}(\pi_{v},\gamma^{*})\leq\epsilon$ with probability at least $1-p$”, where $\epsilon$ and $p$ are certain deterministic, explicit functions of $d$, $n$, and the model. We are only aware of one other work which proves BvMs in this style [26]. Our nonasymptotic BvM results then easily lead to the more classical asymptotic statements.
Before proceeding to our proof outline, we make a few more remarks. First, proving the MLE exists and is unique with high probability is in some sense a prerequisite to proving the BvM. Secondly, for regression models with random design, the random design adds a further source of randomness. Namely, suppose we observe pairs $(X_{i},Y_{i})$, and we have
$p(Y_{1:n}|\theta)=\prod_{i=1}^{n}p(Y_{i}|\theta,X_{i}).$
The $X_{i}$ are drawn from some known distribution $f$ (not modeled in the posterior). Note that the TV distance $\mathrm{TV}(\pi_{v},\gamma^{*})$ is then a random variable whose law depends both on the conditional distribution $Y_{1:n}|X_{1:n}\sim p(\cdot\mid{\theta^{*}})$, and on the design distribution $X_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}f$. In our study of logistic regression with Gaussian design, we will take into account both sources of randomness in our statement of the BvM. Ours seems to be the first BvM to take this extra layer of randomness into account instead of considering the design to be fixed.
3.3 Proof outline and preliminary bounds
As mentioned in the introduction, proving the BvM is similar to proving a Laplace approximation bound with high probability over different draws of the data. We can therefore take advantage of Theorem 2.2 in our BvM proof. Our three main proof steps are as follows. First, we show that discarding the prior has negligible effect by bounding $\mathrm{TV}(\pi_{v},\pi_{\ell})$. Here, $\pi_{\ell}\propto e^{-n\ell}$. Second, we bound $\mathrm{TV}(\pi_{\ell},\gamma_{\ell})$ by directly applying Theorem 2.2 with $f=\ell$. Here, $\gamma_{\ell}=\mathcal{N}(\hat{\theta},(n\nabla^{2}\ell(\hat{\theta}))^{-1})$ is the Laplace approximation to $\pi_{\ell}$. Third, we bound $\mathrm{TV}(\gamma_{\ell},\gamma^{*})$ by comparing the inverse covariance matrices $n\nabla^{2}\ell(\hat{\theta})$ and $n\nabla^{2}\ell_{\infty}({\theta^{*}})$ of the two Gaussian distributions (which have the same mean). We then use the triangle inequality to bound $\mathrm{TV}(\pi_{v},\gamma^{*})$ via the sum of the three bounds.
Together, these three steps give us a random bound on $\mathrm{TV}(\pi_{v},\gamma^{*})$, depending on the random MLE $\hat{\theta}$ and the random function $\ell(\theta)$. To complete the BvM proof, we show that on a high probability event, the point $\hat{\theta}$ lies in a small neighborhood of ${\theta^{*}}$, the derivatives of $\ell$ can be bounded deterministically, and $\nabla^{2}\ell$ is close to $\nabla^{2}\ell_{\infty}$.
We carry out the final probabilistic step for each of our three models separately. But the random bound on $\mathrm{TV}(\pi_{v},\gamma^{*})$ can be stated in a general setting, and we do so in the present section. First we make the following assumption which will allow us to apply Theorem 2.2 to bound $\mathrm{TV}(\pi_{\ell},\gamma_{\ell})$.
Assumption A1.
With probability 1, the negative normalized log likelihood $\ell$ is convex and $C^{3}$ on $\Theta$.
Remark 3.2.
To apply Theorem 2.2, we do not actually need $\ell\in C^{3}(\Theta)$. But from the definition (2.1) of $\delta_{3}(r)$, we see that we (nearly) need $\ell$ to be three times differentiable at the point $\hat{\theta}$. Since now $\hat{\theta}$ is a random point (as it minimizes the random function $\ell$), it is more convenient to simply assume $\ell$ is everywhere three times differentiable.
Throughout the rest of the section, we implicitly assume A1. Next, we introduce several quantities which play an important role in our BvM results.
Definition 3.3.
Let ${F^{*}}$ be a symmetric positive definite matrix and ${\theta^{*}}$ be the ground truth vector. Then we let
$\begin{gathered}{\mathcal{U}}^{*}(s)=\{|\theta-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}\},\qquad\delta^{*}_{3}(s)=\sup_{\theta\in{\mathcal{U}}^{*}(s)}\|\nabla^{3}\ell(\theta)\|_{F^{*}},\\ \delta^{*}_{01}(s)=\sup_{\theta\in{\mathcal{U}}^{*}(s)}\|\nabla\log\pi_{0}(\theta)\|_{F^{*}},\qquad{M_{0}}^{*}=d^{-1}\log\sup_{\theta\in\Theta}\pi_{0}(\theta)/\pi_{0}({\theta^{*}}),\end{gathered}$ (3.2)
with the convention that $\delta^{*}_{01}(s)=\infty$ if the prior is not $C^{1}$ or not strictly positive on ${\mathcal{U}}^{*}(s)$.
Note that ${\mathcal{U}}^{*}(s)$ and $\delta^{*}_{3}(s)$ are analogous to ${\mathcal{U}}(s)$ and $\delta_{3}(s)$ — provided we define the latter to be the righthand side of (2.3) rather than (2.1). The difference between the starred and unstarred quantities is that the former are centered at the point ${\theta^{*}}$ and involve a different weighting matrix ${F^{*}}$. The reason for considering a region centered at ${\theta^{*}}$ rather than $\hat{\theta}$ is that ${\theta^{*}}$ is a deterministic point, whereas $\hat{\theta}$ may not even be defined with some positive probability. Similarly, we also choose ${F^{*}}$ to be a deterministic matrix. In Lemma B.2, we show that if $\hat{\theta}$ is close to ${\theta^{*}}$ then ${\mathcal{U}}(r)$ is a subset of ${\mathcal{U}}^{*}(s)$ and $\delta_{3}(r)\lesssim\delta^{*}_{3}(2s)$ for an $s$ slightly larger than $r$.
Remark 3.4 (Choice of ${F^{*}}$).
The typical choice of ${F^{*}}$ in BvM proofs is ${F^{*}}=\nabla^{2}\ell_{\infty}({\theta^{*}})$, the Fisher information matrix. But for the logistic regression model in Section 6, the Fisher information matrix $\nabla^{2}\ell_{\infty}({\theta^{*}})$ is random due to the random design. To avoid dealing with quantities weighted by a random matrix, in Section 6 we choose a different ${F^{*}}$, namely, the covariance matrix of the feature vector distribution.
The following is the key lemma in the BvM proof. It bounds each of the three TV distances $\mathrm{TV}(\pi_{v},\pi_{\ell}),\mathrm{TV}(\pi_{\ell},\gamma_{\ell}),\mathrm{TV}(\gamma_{\ell},\gamma^{*})$ on a certain random event $E(s,\lambda)$. Note that “$\exists!\,\mathrm{MLE}\,\hat{\theta}$” is an abbreviation for “the MLE exists and is unique”.
Lemma 3.5 (Key preliminary lemma).
Let $s\geq 0,\lambda\in(0,1]$ be such that $s\sqrt{\lambda/2}\geq 6$. Suppose ${\mathcal{U}}^{*}(2s)\subset\Theta$ and $\delta^{*}_{01}(2s)/\sqrt{nd}\leq s\lambda/48$. Then on the event
$E(s,\lambda)=\{\exists!\,\mathrm{MLE}\,\hat{\theta},\,|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}\}\cap\{\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}\}\cap\{2s\delta^{*}_{3}(2s)\sqrt{d/n}\leq\lambda/2\},$ (3.3)
it holds
$\mathrm{TV}(\pi_{\ell},\gamma_{\ell})$ $\leq\frac{2\delta^{*}_{3}(2s)}{\lambda^{3/2}}\,\frac{d}{\sqrt{n}}+3e^{-ds^{2}\lambda/18},$ (3.4)
$\mathrm{TV}(\pi_{v},\pi_{\ell})$ $\leq\frac{\delta^{*}_{01}(2s)}{\sqrt{2\lambda}}\,\frac{1}{\sqrt{n}}+4e^{d[{M_{0}}^{*}-s^{2}\lambda/36]}.$ (3.5)
Furthermore, if $\nabla^{2}\ell_{\infty}({\theta^{*}})\succeq\lambda_{\infty}{F^{*}}$ for some $\lambda_{\infty}>0$, then the following also holds on $E(s,\lambda)$.
$\mathrm{TV}(\gamma_{\ell},\gamma^{*})\leq(\lambda^{-1}+\lambda_{\infty}^{-1})\left(s\delta^{*}_{3}(2s)\frac{d}{\sqrt{n}}+\sqrt{d}\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}\right)$ (3.6)
Remark 3.6.
As noted in Remark 3.4, the typical choice of ${F^{*}}$ is ${F^{*}}=\nabla^{2}\ell_{\infty}({\theta^{*}})$. In this case, $\nabla^{2}\ell_{\infty}({\theta^{*}})\succeq\lambda_{\infty}{F^{*}}$ trivially holds with $\lambda_{\infty}=1$.
The bound (3.4) is a direct application of Theorem 2.2. The proof of (3.5) is similar to the proof of Theorem 2.2, with Lemma 2.9 as the key tool. The bound (3.6) is straightforward since $\gamma_{\ell}$ and $\gamma^{*}$ are Gaussians with the same mean. See Appendix B for the full proof.
Given Lemma 3.5, we can finish the BvM proof by 1) bounding from below the probability of $E(s,\lambda)$ and 2) bounding $\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}$ and $\delta^{*}_{3}(2s)$ with high probability (w.h.p.). The second step is needed to obtain further, deterministic bounds on the righthand sides of (3.4) and (3.6). Note that 2) is not really separate from 1). Indeed,
$\mathbb{P}\left(\delta^{*}_{3}(2s)\leq c_{3}\right)\;\leq\;\mathbb{P}\left(2s\delta^{*}_{3}(2s)\sqrt{d/n}\leq\lambda/2\right)\quad\forall\;\lambda\geq 4sc_{3}\sqrt{d/n}.$
Also, in the most common case that ${F^{*}}=\nabla^{2}\ell_{\infty}({\theta^{*}})$, we have that
$\mathbb{P}\left(\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}\leq\epsilon\right)\;\leq\;\mathbb{P}\left(\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}\right)\quad\forall\;\lambda\leq 1-\epsilon.$
Therefore, bounding $\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}$ and $\delta^{*}_{3}(2s)$ w.h.p. proves the second and third events in $E(s,\lambda)$ hold w.h.p.
Next, we provide conditions which ensure the first event holds, i.e. that there exists a unique MLE $\hat{\theta}$, and that this point is close to ${\theta^{*}}$. In two of our three example applications (exponential families and logistic regression), it will be simpler to bound the probability that the sufficient condition holds than to directly bound the probability that a unique MLE close to ${\theta^{*}}$ exists.
Lemma 3.7.
Suppose ${\mathcal{U}}^{*}(s)\subset\Theta$ for some $s\geq 0$, and let $0<\lambda\leq 1$. Then on the event
$\bar{E}(s,\lambda)=\{\|\nabla\ell({\theta^{*}})\|_{F^{*}}\leq 2\lambda s\sqrt{d/n}\}\cap\{\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}\}\cap\{2s\delta^{*}_{3}(2s)\sqrt{d/n}\leq\lambda/2\}$ (3.7)
the function $\ell$ has a unique global minimizer $\hat{\theta}$, and $|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}$. Hence $\bar{E}(s,\lambda)\subset E(s,\lambda)$.
See Appendix B for the proof.
Prior assumptions.
We conclude this section by discussing the prior. Firstly, note that the bound (3.5) from the prior contribution is already deterministic. Thus in any statistical model for which $E(s,\lambda)$ occurs with high probability for an absolute constant $\lambda$ bounded away from zero,  (3.5) implies
[…]
Formally, our only assumption on the prior is that $\delta^{*}_{01}(2s)/\sqrt{nd}\leq s\lambda/48$, which for bounded $s$ reads $\delta^{*}_{01}(2s)=\mathcal{O}(\sqrt{nd})$. Practically, however, the bound (3.8) itself imposes the stronger assumption that
${M_{0}}^{*}=\mathcal{O}(1),\qquad\delta^{*}_{01}(s)=o(\sqrt{n}).$ (3.9)
Before moving on to compare our prior assumptions (3.9) to those of past works, let us compute $\delta^{*}_{01}$ and ${M_{0}}^{*}$ for a few choices of prior.
Example 3.8.
Flat prior: $\pi_{0}(\theta)\equiv 1$. Then ${M_{0}}^{*}=0$, $\delta^{*}_{01}(s)\equiv 0.$
Gaussian prior $\pi_{0}=\mathcal{N}(\mu,\Sigma)$:
$\begin{split}{M_{0}}^{*}&\leq\frac{1}{2d}\|\Sigma^{-1}\|_{{F^{*}}}|\mu-{\theta^{*}}|_{F^{*}}^{2},\\ \delta^{*}_{01}(s)&\leq\|\Sigma^{-1}\|_{F^{*}}(|\mu-{\theta^{*}}|_{F^{*}}+s\sqrt{d/n}).\end{split}$ (3.10)
Multivariate Student’s t prior $\pi_{0}=t_{\nu}(\mu,\Sigma)$:
$\begin{split}{M_{0}}^{*}&\leq\frac{\nu+d}{2\nu d}\|\Sigma^{-1}\|_{F^{*}}|\mu-{\theta^{*}}|_{F^{*}}^{2},\\ \delta^{*}_{01}(s)&\leq\frac{\nu+d}{\nu}\|\Sigma^{-1}\|_{F^{*}}(|\mu-{\theta^{*}}|_{F^{*}}+s\sqrt{d/n}).\end{split}$ (3.11)
For specificity, consider now the case in which $s\sqrt{d/n}\leq 1$, $|\mu-{\theta^{*}}|_{F^{*}}=1$, and $\nu=\mathcal{O}(1)$ in the Student’s t prior. Suppose $\Sigma^{-1}=w{F^{*}}$ so that larger $w$ (for “weight”) corresponds to stronger regularization. Indeed, the covariance of both the Gaussian and Student’s t prior are proportional to $\Sigma=w^{-1}{F^{*}}^{-1}$, so larger $w$ corresponds to a stronger belief that the parameter $\theta$ is near the mean $\mu$. We then obtain the following bounds:
$\begin{split}\text{Gaussian prior:}\qquad&{M_{0}}^{*}\lesssim d^{-1}w,\qquad\delta^{*}_{01}(s)\lesssim w,\\ \text{Student's t prior}:\qquad&{M_{0}}^{*}\lesssim w,\qquad\delta^{*}_{01}(s)\lesssim dw\end{split}$ (3.12)
See Appendix B for the calculations in this example. The condition ${M_{0}}^{*}=\mathcal{O}(1)$ is standard and has been assumed in BvM proofs in the works [9, 10, 5, 18, 26]. However, we have assumed slightly stronger regularity of the log prior near ${\theta^{*}}$. Namely, we require that $\log\pi_{0}\in C^{1}({\mathcal{U}}^{*}(s))$, while the above five works only require that $\log\pi_{0}$ has a sufficiently small Lipschitz constant near ${\theta^{*}}$. For the sake of comparison, let us denote this Lipschitz constant $\delta^{*}_{01}(s)$ also, since it follows from a bound on the first derivative. For the small price of strengthening the regularity from Lipschitz to $C^{1}$, we can use a log Sobolev inequality in ${\mathcal{U}}^{*}(s)$ to obtain a tighter bound on $\mathrm{TV}(\pi_{\ell},\pi_{v})$, showing that $\delta^{*}_{01}(s)=o(\sqrt{n})$ suffices for this TV distance to be negligible. In contrast, the BvMs in the above five works require $\delta^{*}_{01}(s)=o(\sqrt{n/d})$. Similarly, one can show that contribution from the prior in the “MLE-centered” Laplace approximation bound of [13] is negligible only if $\delta^{*}_{01}(s)=\mathcal{O}(\sqrt{n/d})$.
We have therefore shown the strength of prior regularization can be greater than was previously known. To illustrate this for the Student’s t prior in Example 3.8, the regularization strength parameter $w$ can be as large as $w\sim\sqrt{n}/d$, in contrast to $w\sim\sqrt{n}/d\sqrt{d}$ previously.
4 Application to Exponential Families
In this section, we prove the BvM for exponential families in increasing dimension. We describe the set-up in Section 4.1, prove the BvM in Section 4.2, and apply the result to the special case of log concave exponential families in Section 4.3. The exponential family setting was first considered in [10], and revisited in [5]. We compare our result to these two works at the end of Section 4.2.
4.1 Set-up
In the exponential family setup, we observe i.i.d. data $Y_{i}\in\mathbb{R}^{d}$, $i=1,\dots,n$, modeled as $Y_{i}\sim p(\cdot\mid\theta)$ for some unknown $\theta\in\Theta\subset\mathbb{R}^{d}$. Here, $p$ is an exponential family of the form
$p(y\mid\theta)=h(y)\,\mathrm{exp}\left(\theta^{\intercal}y-\psi(\theta)\right),$ (4.1)
for some base density $h$. The derivatives of $\psi$ can be expressed in terms of moments of $p(\cdot\mid\theta)$:
$\begin{split}\nabla\psi(\theta)&=\mathbb{E}\,_{\theta}[Y],\\ \nabla^{2}\psi(\theta)&={\mathrm{Var}}_{\theta}(Y),\\ \nabla^{3}\psi(\theta)&=\mathbb{E}\,_{\theta}\left[(Y-\mathbb{E}\,_{\theta}[Y])^{\otimes 3}\right],\end{split}$ (4.2)
where $\mathbb{E}\,_{\theta}[Y]$ is shorthand for $\mathbb{E}\,_{Y\sim p(\cdot\mid\theta)}[Y]$. This follows from the fact that $\partial_{\theta_{1}}^{k_{1}}\dots\partial_{\theta_{d}}^{k_{d}}\psi(\theta)$ is the order $(k_{1},\dots,k_{d})$ cumulant of $p(\cdot\mid\theta)$; see Theorem 8.1 in [4].
Now, the model (4.1) leads to the following normalized negative log likelihood and negative population log likelihood:
$\begin{split}\ell(\theta)&=-\frac{1}{n}\sum_{i=1}^{n}\log p(y_{i}\mid\theta)=\psi(\theta)-\theta^{\intercal}\bar{Y},\\ \ell_{\infty}(\theta)&=\mathbb{E}\,[\ell(\theta)]=\psi(\theta)-\theta^{\intercal}\mathbb{E}\,_{{\theta^{*}}}[Y].\end{split}$ (4.3)
where $\bar{Y}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}$. Taking the derivatives of $\ell$ and using (4.2) gives
$\nabla\ell(\theta)$ $=\nabla\psi(\theta)-\bar{Y}=\mathbb{E}\,_{\theta}[Y]-\bar{Y}$ (4.4)
$\nabla^{2}\ell(\theta)$ $=\nabla^{2}\psi(\theta)={\mathrm{Var}}_{\theta}(Y),$ (4.5)
$\nabla^{3}\ell(\theta)$ $=\nabla^{3}\psi(\theta)=\mathbb{E}\,_{\theta}\left[(Y-\mathbb{E}\,_{\theta}[Y])^{\otimes 3}\right].$ (4.6)
Note that (4.5) shows that $\ell$ is always convex, since ${\mathrm{Var}}_{\theta}(Y)\succeq 0$. Also, from (4.6) we see that $\nabla^{3}\ell\in C^{3}(\Theta)$ when $\mathbb{E}\,_{\theta}[|Y|^{3}]$ is bounded uniformly over compact subsets of $\Theta$. In this case, A1 is satisfied. Furthermore, we see that $\nabla^{2}\ell({\theta^{*}})=\nabla^{2}\ell_{\infty}({\theta^{*}})$ is satisfied for this model.
For the matrix ${F^{*}}$, we choose ${F^{*}}=\nabla^{2}\ell({\theta^{*}})=\nabla^{2}\ell_{\infty}({\theta^{*}})={\mathrm{Var}}_{{\theta^{*}}}(Y)$, the Fisher information. Therefore ${\mathcal{U}}^{*}(t)$, $\delta^{*}_{3}$, and $\delta^{*}_{01}$ from Definition 3.3 take the form
$\begin{split}{\mathcal{U}}^{*}(t)=&\left\{\theta\;:\;{\mathrm{Var}}_{{\theta^{*}}}((\theta-{\theta^{*}})^{\intercal}Y)\leq t^{2}d/n\right\},\\ \delta^{*}_{3}(t)=&\sup_{\theta\in{\mathcal{U}}^{*}(t)}\;\sup_{u:{\mathrm{Var}}_{{\theta^{*}}}(u^{\intercal}Y)=1}\mathbb{E}\,_{\theta}\left[(u^{\intercal}Y-\mathbb{E}\,_{\theta}[u^{\intercal}Y])^{3}\right],\\ \delta^{*}_{01}(t)=&\sup_{\theta\in{\mathcal{U}}^{*}(t)}\;\sup_{u:{\mathrm{Var}}_{{\theta^{*}}}(u^{\intercal}Y)=1}u^{\intercal}\nabla\log\pi_{0}(\theta).\end{split}$ (4.7)
4.2 BvM proof
To prove the BvM we use Lemma 3.5 together with Lemma 3.7, which shows that $\bar{E}(s,\lambda)$ from the latter lemma is a subset of $E(s,\lambda)$ from the former lemma. Thus it suffices to bound the probability of $\bar{E}(s,\lambda)$. Note that the usual additional step of bounding $\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}$ and $\delta^{*}_{3}(2s)$ with high probability (to get deterministic righthand sides in (3.4) and (3.6)) is moot, since $\nabla^{2}\ell_{\infty}=\nabla^{2}\ell$ and $\delta^{*}_{3}$ is deterministic. Consider the definition (3.7) of $\bar{E}(s,\lambda)$ for $\lambda=1$. We see that the second event is trivially satisfied with probability 1, since $\nabla^{2}\ell({\theta^{*}})={F^{*}}$. Moreover, the third event also drops out, since it becomes a condition on the deterministic $\delta^{*}_{3}$. Thus $\bar{E}(s,1)$ reduces to the first event, i.e. $\bar{E}(s,1)=\{\|\nabla\ell({\theta^{*}})\|_{F^{*}}\leq 2s\sqrt{d/n}\}$. We now bound the probability of this event from below.
Lemma 4.1.
Let $s\geq 0$ be such that ${\mathcal{U}}^{*}(s)\subset\Theta$ and $s\delta^{*}_{3}(s)\sqrt{d/n}\leq 1/4$. Then
$\mathbb{P}\left(\left|{\mathrm{Var}}_{\theta^{*}}(Y)^{-1/2}(\bar{Y}-\mathbb{E}\,_{\theta^{*}}[Y])\right|\leq 2s\sqrt{d/n}\right)\geq 1-\,\mathrm{exp}(d[\log 5-s^{2}/4]).$
under the distribution $Y_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}p(\cdot\mid{\theta^{*}})$, $i=1,\dots,n$.
Note that $|{\mathrm{Var}}_{\theta^{*}}(Y)^{-1/2}(\bar{Y}-\mathbb{E}\,_{\theta^{*}}[Y])|$ is precisely $\|\nabla\ell({\theta^{*}})\|_{F^{*}}$. If we multiply this quantity by $\sqrt{n}$, we get the norm of a standardized sample average. By the Central Limit Theorem, we expect this quantity to behave like the norm of a $d$-dimensional standard Gaussian vector. Indeed, the tail probability $e^{d(\log 5-s^{2}/4)}$ in the lemma is like that of a standard Gaussian (up to suboptimal constants). The asymptotic distribution of the norm of a standardized sample mean was first considered in [22]. Later, [28, Section F.3] proved nonasymptotic tail bounds on the norm of a sub-Gaussian or sub-exponential random vector. The results in this latter work could be applied to our setting, but to be self-contained we give our own proof of Lemma 4.1 using the specifics of our setting; see Appendix C.
We now apply Lemma 3.5 with $\lambda=\lambda_{\infty}=1$ to deduce the BvM.
Proposition 4.2 (Non-asymptotic BvM for exponential families).
Let $s\geq 6\sqrt{2}$, and suppose ${\mathcal{U}}^{*}(2s)\subset\Theta$, $s\delta^{*}_{3}(2s)\sqrt{d/n}\leq 1/4$, and $\delta^{*}_{01}(2s)\leq s\sqrt{nd}/48$, where ${\mathcal{U}}^{*}$, $\delta^{*}_{3}$, $\delta^{*}_{01}$ are as in (4.7). Then on an event of probability at least $1-e^{-ds^{2}/8}$ there exists a unique MLE $\hat{\theta}$, and it holds
$\begin{split}|\hat{\theta}-{\theta^{*}}|_{F^{*}}&\leq s\sqrt{d/n},\\ \mathrm{TV}(\pi_{v},\pi_{\ell})&\leq\frac{\delta^{*}_{01}(2s)}{\sqrt{2n}}+4e^{d({M_{0}}^{*}-s^{2}/36)},\\ \mathrm{TV}(\pi_{\ell},\gamma_{\ell})&\leq 2\delta^{*}_{3}(2s)\frac{d}{\sqrt{n}}+3e^{-ds^{2}/18},\\ \mathrm{TV}\left(\gamma_{\ell},\gamma^{*}\right)&\leq 2s\delta^{*}_{3}(s)\frac{d}{\sqrt{n}}.\end{split}$ (4.8)
The proof follows directly from Lemmas 3.5 and 4.1. Note also that $e^{-d(\log 5-s^{2}/4)}\leq e^{-ds^{2}/8}$ when $s\geq 6\sqrt{2}$. We now prove the traditional asymptotic BvM by adding the three TV bounds from Proposition 4.2. We then take the $n\to\infty$ limit for each fixed $s$, and finally take $s\to\infty$.
Corollary 4.3 (Asymptotic BvM for exponential families).
Let $d=d_{n}$ and $\Theta=\Theta_{n}$. Similarly, write $v_{n},\gamma^{*}_{n},{\mathcal{U}}^{*}_{n},\delta^{*}_{3n},M_{0n}^{*},\delta^{*}_{01n}$ to emphasize the dependence of these quantities on $n$. Suppose $M_{0n}^{*}=\mathcal{O}(1)$ as $n\to\infty$ and that for each $s\geq 0$, the following hold: (1) the neighborhood ${\mathcal{U}}^{*}_{n}(s)$ is contained in $\Theta_{n}$ when $n$ is large enough, (2) $\delta^{*}_{01n}(s)=o(\sqrt{n})$, and (3)
$\delta^{*}_{3n}(s)\frac{d_{n}}{\sqrt{n}}=o(1).$
Then $\mathrm{TV}(\pi_{v_{n}},\gamma^{*}_{n})\to 0$ with probability tending to 1 as $n\to\infty$.
Let us compare Corollary 4.3 to the exponential family BvMs in [10] and [5]. We have already compared our conditions on the prior to those in these two works in Section 3.3. The most substantial difference in the BvM results is that we require only that $\delta^{*}_{3n}(s)d_{n}/\sqrt{n}=o(1)$, whereas [5] assumes $\delta^{*}_{3n}(0)d_{n}\sqrt{d_{n}}/\sqrt{n}=o(1)$ as well as an additional growth condition on the fourth derivative. (Combined, the third and fourth derivative assumptions imply $\delta^{*}_{3n}(s)d_{n}\sqrt{d_{n}}/\sqrt{n}=o(1)$.) The work [5] in turn improves on that of [10] by removing a $\log d_{n}$ factor.
4.3 Log concave data distribution
Similarly to [5], we now consider the case in which the exponential family $p(y|\theta)$ from (4.1) is log concave in $y$. The statement of the BvM simplifies in this case because the norm of the third moment tensor can be bounded by the norm of the matrix ${\mathrm{Var}}_{\theta}(Y)$.
Lemma 4.4.
Suppose the base density $h$ from (4.1) is log concave in $y$. Let
$\delta^{*}_{2}(s)=\sup_{\theta\in{\mathcal{U}}^{*}(s)}\|F({\theta^{*}})^{-1/2}F(\theta)F({\theta^{*}})^{-1/2}\|=\sup_{\theta\in{\mathcal{U}}^{*}(s)}\|F(\theta)\|_{F^{*}},$
where $F(\theta)={\mathrm{Var}}_{\theta}(Y)$. Then $\delta^{*}_{3}(s)\leq C_{23}\delta^{*}_{2}(s)^{3/2}$ for some absolute constant $C_{23}$.
Proof.
If $h$ is log concave then $p(\cdot|\theta)$ is log concave for each $\theta\in\Theta$. Now, log concavity is preserved under affine transformations  [23, Section 3.1.1]. Therefore if $Y\sim p(\cdot\mid\theta)$ then $u^{\intercal}Y-\mathbb{E}\,_{\theta}[u^{\intercal}Y]$ also has a log concave distribution for all $u\in\mathbb{R}^{d}$. [17, Theorem 5.22] now gives that
$\mathbb{E}\,_{\theta}\left[\left|u^{\intercal}Y-\mathbb{E}\,_{\theta}[u^{\intercal}Y]\right|^{3}\right]\leq C_{23}{\mathrm{Var}}_{\theta}(u^{\intercal}Y)^{3/2}=C_{23}\left(u^{\intercal}F(\theta)u\right)^{3/2}.$
Taking the supremum over all $u$ such that $|u|_{F^{*}}^{2}=1$ and $\theta\in{\mathcal{U}}^{*}(s)$ concludes the proof.∎
Using this result in Proposition 4.2 leads to the following simplified BvM. We do not indicate the bound on $\mathrm{TV}(\pi_{v},\pi_{\ell})$, since it is not affected by log concavity of $p(\cdot\mid\theta)$.
Proposition 4.5 (Nonasymptotic BvM for log concave exponential families).
Suppose the base density $h$ of the exponential family (4.1) is log concave. Suppose $s\geq 6\sqrt{2}$, $s\sqrt{d/n}\leq(8C_{23})^{-1}$, ${\mathcal{U}}^{*}(2s)\subset\Theta$, and
$\delta^{*}_{2}(2s)=\sup_{\theta\in{\mathcal{U}}^{*}(2s)}\|F({\theta^{*}})^{-1/2}F(\theta)F({\theta^{*}})^{-1/2}\|\leq\frac{3}{2},$ (4.9)
where $F(\theta):={\mathrm{Var}}_{\theta}(Y)$. Then on an event of probability at least $1-e^{-ds^{2}/8}$ there exists a unique MLE $\hat{\theta}$ with $|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}$, and it holds
$\begin{split}\mathrm{TV}\left(\gamma_{\ell},\gamma^{*}\right)\leq Cs\frac{d}{\sqrt{n}},\qquad\mathrm{TV}(\pi_{\ell},\gamma_{\ell})\leq C\frac{d}{\sqrt{n}}+3e^{-ds^{2}/18}.\end{split}$ (4.10)
Proof.
The assumptions and Lemma 4.4 give $s\delta^{*}_{3}(2s)\sqrt{d/n}\leq(3/2)^{3/2}C_{23}s\sqrt{d/n}\leq 2C_{23}s\sqrt{d/n}\leq 1/4.$ Therefore the assumptions of Proposition 4.2 are satisfied. Using the bound $\delta^{*}_{3}(2s)\leq(3/2)^{3/2}C_{23}$ in the righthand sides of (4.8) concludes the proof of (4.10). ∎
A nice feature of Proposition 4.5 is that it only requires one to check (4.9), which is a condition involving matrix norms alone; no 3-tensor norms need to be computed. Moreover, the bounds (4.10) are in general stronger than the corresponding bounds in Proposition 4.2, because the coefficients in (4.10) are always independent of $d$. In contrast, the coefficient $\delta^{*}_{3}(2s)$ in Proposition 4.2 can depend on $d$. For example, in the discrete distribution model in Section 5 (which has an exponential family parameterization), Remark 5.9 shows that even at $s=0$, we have $\delta^{*}_{3}(0)\gtrsim\sqrt{d}$.
Of course, the reason the coefficients in (4.10) are independent of $d$ is that we simply assumed $\delta^{*}_{2}(2s)\leq 3/2$. This may seem like a strong assumption, but actually it is not. To see why, first note that $\delta^{*}_{2}(0)=\|{F^{*}}\|_{F^{*}}=1$. Therefore when $d/n$ is small, we can expect $\delta^{*}_{2}(2s)$ to be bounded. To be more precise, we show that $\delta^{*}_{2}(2s)\leq 3/2$ is implied by (and hence weaker than) the assumption $s\delta^{*}_{3}(2s)\sqrt{d/n}\leq 1/4$ from Proposition 4.2. Indeed, Taylor’s theorem gives
$\delta^{*}_{2}(2s)=\sup_{\theta\in{\mathcal{U}}^{*}(2s)}\|\nabla^{2}\ell(\theta)\|_{F^{*}}\leq 1+\sup_{\theta\in{\mathcal{U}}^{*}(2s)}\|\nabla^{2}\ell(\theta)-\nabla^{2}\ell({\theta^{*}})\|_{F^{*}}\leq 1+\delta^{*}_{3}(2s)2s\sqrt{d/n}\leq\frac{3}{2}.$
Thus, log concavity alone explains why the coefficients in (4.10) are independent of $d$.
5 Observations of a discrete probability distribution
In this section, we consider a model in which the parameter of interest are the probabilities of a finite state probability mass function (pmf) ${\theta^{*}}=(\theta_{0}^{*},\theta_{1}^{*},\dots,\theta_{d}^{*})$, and we are given $n$ i.i.d. draws from ${\theta^{*}}$. This is equivalent to observing a single sample $Y\sim\mathrm{Multi}(n,{\theta^{*}})$, the multinomial distribution with $n$ trials and probabilities $\theta_{0}^{*},\theta_{1}^{*},\dots,\theta_{d}^{*}$. A BvM for this model, in which $d$ can grow large, was first proved in [6]. We compare our result to that of [6] at the end of Section 5.2.
There is a reparameterization which turns the model into an exponential family; the works [10] and [5] on BvMs for exponential families both studied it in this form. However, a BvM for one form of the model does not immediately imply the BvM for the other form; as noted in [6], the effect of reparameterization on asymptotic normality must first be established.
Proofs of omitted results in this section can be found in Appendix D.
5.1 Set-up
The $d+1$ probabilities of a pmf over states $0,1,\dots,d$ add up to 1. We therefore define the parameter space to be the probabilities of states $1,\dots,d$ only:
$\begin{split}\Theta=\{\theta=(\theta_{1},\dots,\theta_{d})\mid 0<\theta_{1}+\dots+\theta_{d}<1\}.\end{split}$ (5.1)
The cases $\theta_{1}+\dots+\theta_{d}=0$ and $\theta_{1}+\dots+\theta_{d}=1$ are degenerate, so we exclude them. Now, to every point $\theta\in\Theta$, we associate the value
$\theta_{0}=1-\sum_{j=1}^{d}\theta_{j}=1-\mathds{1}^{\intercal}\theta.$
Note that if $\theta\in\Theta$ then $(\theta_{0},\dots,\theta_{d})$ is a pmf on $d+1$ states such that each probability $\theta_{j},j=0,\dots,d$ is strictly positive. We will often abuse notation by interchangeably using $\theta$ to denote either $(\theta_{0},\dots,\theta_{d})$ or $(\theta_{1},\dots,\theta_{d})$. One need only remember that $\theta_{1},\dots,\theta_{d}$ are free parameters, while $\theta_{0}$ is determined from the others. We observe counts $N=(N_{0},N_{1},\dots,N_{d})\sim\mathrm{Multi}(n,{\theta^{*}})$, where ${\theta^{*}}=(\theta_{0}^{*},\theta_{1}^{*},\dots,\theta_{d}^{*})$ is the ground truth pmf for some $(\theta_{1}^{*},\dots,\theta_{d}^{*})\in\Theta$. For the proportions corresponding to the counts, we use the notation
$\bar{N}=(\bar{N}_{0},\dots,\bar{N}_{d}),\quad\bar{N}_{j}:=\frac{1}{n}N_{j}.$
Define also
${\theta_{\mathrm{min}}^{*}}=\min_{j=0,\dots,d}\theta_{j}^{*},$
which will play an important role in describing how far ${\theta^{*}}$ is from the boundary of $\Theta$. Now, the multinomial observations give rise to the likelihood $L(\theta)=\prod_{j=0}^{d}\theta_{j}^{N_{j}}$, and negative normalized log likelihood $\ell=-\frac{1}{n}\log L$:
$\ell(\theta)=-\sum_{j=0}^{d}\bar{N}_{j}\log\theta_{j}=-\bar{N}_{0}\log(1-\mathds{1}^{\intercal}\theta)-\sum_{j=1}^{d}\bar{N}_{j}\log\theta_{j}.$ (5.2)
The first three derivatives of $\ell$, with respect to the free parameters $\theta_{1},\dots,\theta_{d}$, are given as follows:
$\begin{split}\nabla\ell(\theta)&=-\bigg{(}\frac{\bar{N}_{j}}{\theta_{j}}\bigg{)}_{j=1}^{d}+\frac{\bar{N}_{0}}{\theta_{0}}\mathds{1},\\ \nabla^{2}\ell(\theta)&=\mathrm{diag}\bigg{(}\bigg{(}\frac{\bar{N}_{j}}{\theta_{j}^{2}}\bigg{)}_{j=1}^{d}\bigg{)}+\frac{\bar{N}_{0}}{\theta_{0}^{2}}\mathds{1}\mathds{1}^{\intercal}\\ \nabla^{3}\ell(\theta)&=-2\mathrm{diag}\bigg{(}\bigg{(}\frac{\bar{N}_{j}}{\theta_{j}^{3}}\bigg{)}_{j=1}^{d}\bigg{)}+2\frac{\bar{N}_{0}}{\theta_{0}^{3}}\mathds{1}^{\otimes 3}\end{split}$ (5.3)
Remark 5.1.
We see from the second equation that $\nabla^{2}\ell(\theta)\succeq 0$ for all $\theta\in\Theta$ with probability 1. Hence, $\ell$ is convex on $\Theta$ and clearly $\ell\in C^{3}(\Theta)$, so Assumption A1 is satisfied.
Remark 5.2.
Suppose $\bar{N}\in\Theta$, which implies the counts $N_{j}$, $j=0,\dots,d$ are all strictly positive. Since $\ell$ is convex, any strict local minimizer $\hat{\theta}$ of $\ell$ in $\Theta$ must be the unique global minimizer, i.e. the MLE. But we see from the first two equations that $\nabla\ell(\bar{N})=0$ and $\nabla^{2}\ell(\bar{N})\succ 0$. Thus $\bar{N}$ is the unique MLE as long as $\bar{N}\in\Theta$.
For the weighting matrix ${F^{*}}$, we use the Fisher information matrix ${F^{*}}=\nabla^{2}\ell_{\infty}({\theta^{*}})=\mathbb{E}\,[\nabla^{2}\ell({\theta^{*}})]$. Using that $\mathbb{E}\,[\bar{N}]={\theta^{*}}$ gives
${F^{*}}=\nabla^{2}\ell_{\infty}({\theta^{*}})=\mathrm{diag}\bigg{(}\bigg{(}\frac{1}{\theta_{j}^{*}}\bigg{)}_{j=1}^{d}\bigg{)}+\frac{1}{\theta_{0}^{*}}\mathds{1}\mathds{1}^{\intercal}$ (5.4)
As usual, we define
${\mathcal{U}}^{*}(r)=\{\theta\in\mathbb{R}^{d}:|\theta-{\theta^{*}}|_{F^{*}}\leq r\sqrt{d/n}\}.$ (5.5)
Definition 5.3 (Chi-squared divergence).
Let $\theta=(\theta_{0},\dots,\theta_{d})$ be a strictly positive pmf on $d+1$ states. Let $\omega$ be either $\omega=(\omega_{0},\dots,\omega_{d})\in\mathbb{R}^{d+1}$ such that $\omega^{\intercal}\mathds{1}=1$, or $\omega=(\omega_{1},\dots,\omega_{d})\in\mathbb{R}^{d}$. In the latter case, set $\omega_{0}=1-\sum_{j=1}^{d}\omega_{j}$. Then we define
$\chi^{2}(\omega||\theta)=\sum_{j=0}^{d}(\omega_{j}-\theta_{j})^{2}/\theta_{j}.$
Remark 5.4.
We will make frequent use of the bound
$\max_{j=0,\dots,d}\left|\frac{\theta_{j}}{\theta_{j}^{*}}-1\right|^{2}\leq\chi^{2}(\theta||{\theta^{*}})/{\theta_{\mathrm{min}}^{*}},$ (5.6)
which follow directly from the above definition of $\chi^{2}$ divergence.
We now show that the neighborhoods ${\mathcal{U}}^{*}(r)$ from (5.5) are $\chi^{2}$ balls around ${\theta^{*}}$. We also characterize how large $r$ can be to ensure ${\mathcal{U}}^{*}(r)$ remains a subset of $\Theta$, and describe a useful property of pmfs in ${\mathcal{U}}^{*}(r)$
Lemma 5.5.
The neighborhoods ${\mathcal{U}}^{*}(r)$ from (5.5) are equivalently given by
${\mathcal{U}}^{*}(r)=\{\theta\in\mathbb{R}^{d}\;:\;\chi^{2}(\theta||{\theta^{*}})\leq r^{2}d/n\}.$
If $r^{2}d/n<{\theta_{\mathrm{min}}^{*}}/4$, then ${\mathcal{U}}^{*}(2r)\subset\Theta$ and ${\mathcal{U}}^{*}(r)\subset\{\theta\in\Theta\;:\theta_{j}\geq\theta_{j}^{*}/2\;\,\forall j=0,1,\dots,d\}$.
Remark 5.6.
The lemma shows that how large $r$ can be to ensure ${\mathcal{U}}^{*}(2r)\subset\Theta$ depends on how small ${\theta_{\mathrm{min}}^{*}}$ is. This explains our earlier comment that ${\theta_{\mathrm{min}}^{*}}$ controls how far ${\theta^{*}}$ is from the boundary of $\Theta$.
5.2 BvM proof
In this section, we prove the BvM by applying Lemma 3.5. We show that on the event $E(s):=\{\bar{N}\in{\mathcal{U}}^{*}(s)\}$ the following three statements hold:
(1) there exists a unique MLE $\hat{\theta}$ satisfying $|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}$,
(2) it holds $\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}\leq\sqrt{s^{2}d/n{\theta_{\mathrm{min}}^{*}}}$,
(3) $\delta^{*}_{3}(2s)\lesssim 1/\sqrt{{\theta_{\mathrm{min}}^{*}}}$.
The bound on the difference between Hessians implies a lower bound $\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}$ (since $\nabla^{2}\ell_{\infty}({\theta^{*}})={F^{*}}$), and the bound on $\delta^{*}_{3}(2s)$ implies a bound on $2s\delta^{*}_{3}(2s)\sqrt{d/n}$. In other words, the three statements imply $E(s)\subset E(s,\lambda)$ for some suitable $\lambda$. At the same time, the second and third statements yield deterministic bounds on the righthand sides of (3.4) and (3.6). Finally, $\nabla^{2}\ell_{\infty}({\theta^{*}})\succeq\lambda_{\infty}{F^{*}}$ is trivially satisfied with $\lambda_{\infty}=1$. Therefore, proving the three statements and bounding the probability of $E(s)$ will conclude the BvM proof.
The first statement is immediate if we assume $s^{2}d/n\leq{\theta_{\mathrm{min}}^{*}}/4$. Indeed, this assumption implies ${\mathcal{U}}^{*}(s)\subset\Theta$ by Lemma 5.5, and hence on $E(s)$ we have $\bar{N}\in{\mathcal{U}}^{*}(s)\subset\Theta$. But Remark 5.2 then shows $\hat{\theta}=\bar{N}$ is the unique MLE, and since $\bar{N}\in{\mathcal{U}}^{*}(s)$ we have $|\bar{N}-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}$. In the following lemmas, we bound the probability of $E(s)$ and prove statements (2) and (3).
Lemma 5.7.
There is an absolute constant $C_{1}$ such that if $s\geq C_{1}$ and $s^{2}d/n\leq\min(C_{1}^{2},\frac{1}{4}){\theta_{\mathrm{min}}^{*}}$, then $E(s)=\{\chi^{2}(\bar{N}||{\theta^{*}})\leq s^{2}d/n\}$ has probability at least $1-e^{-s^{2}d/C_{1}^{2}}$.
To prove the above lemma, we use a result from [6]’s BvM proof for this model; the result is in turn based on Talagrand’s inequality for the suprema of empirical processes [20].
Lemma 5.8.
Let $s\geq C_{1}$ and $s^{2}d/n\leq\min(C_{1}^{2},\frac{1}{16}){\theta_{\mathrm{min}}^{*}}$. Then $\delta^{*}_{3}(2s)\leq\frac{C_{3}}{\sqrt{\theta_{\mathrm{min}}^{*}}}$ on $E(s)$ for an absolute constant $C_{3}\geq 1$.
Remark 5.9.
The upper bound $\delta^{*}_{3}(2s)\leq 1/\sqrt{{\theta_{\mathrm{min}}^{*}}}$ is tight. To see this, suppose for simplicity that $\bar{N}_{j}=\theta_{j}^{*}$ exactly. This allows us to compute $\|\nabla^{3}\ell({\theta^{*}})\|_{F^{*}}$ explicitly; see Appendix D. The result is that
$\delta^{*}_{3}(2s)\geq\delta^{*}_{3}(0)=\|\nabla^{3}\ell({\theta^{*}})\|_{F^{*}}=2\frac{1-2{\theta_{\mathrm{min}}^{*}}}{\sqrt{{\theta_{\mathrm{min}}^{*}}}\sqrt{1-{\theta_{\mathrm{min}}^{*}}}}\sim\frac{1}{\sqrt{\theta_{\mathrm{min}}^{*}}}$ (5.7)
Since the ground truth pmf values add up to 1, it follows that ${\theta_{\mathrm{min}}^{*}}\leq 1/(d+1)$. Thus the smallest $\delta^{*}_{3}(2s)$ could possibly be is $\delta^{*}_{3}(2s)\sim\sqrt{d}$.
Lemma 5.10.
Let $s\geq C_{1}$ and $s^{2}d/n\leq\min(C_{1}^{2},\frac{1}{4}){\theta_{\mathrm{min}}^{*}}$. Then on $E(s)$, it holds
$\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}\leq\sqrt{\frac{s^{2}d}{n{\theta_{\mathrm{min}}^{*}}}}.$ (5.8)
We conclude by Lemmas 5.8 and 5.10 that if $\sqrt{\frac{s^{2}d}{n{\theta_{\mathrm{min}}^{*}}}}\leq 1/(5C_{3})\leq 1/5$ then $52s\delta^{*}_{3}(2s)\sqrt{d/n}\leq 2/$ 5 and $\nabla^{2}\ell({\theta^{*}})\succeq\frac{4}{5}{F^{*}}$ on $E(s)$. Combining this with statement (1) above, we conclude that $E(s)\subset E(s,4/5)$. We now apply Lemma 3.5 with $\lambda_{\infty}=1,\lambda=4/5$, substituting the deterministic bounds on $\delta^{*}_{3}(2s)$ and $\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}$ from the above lemmas into the bounds (3.4), (3.6). We obtain the following BvM.
Corollary 5.11.
Suppose $s\geq\max(C_{1},6\sqrt{3})$ and $*s^{2}d/n\leq\min(C_{1},\frac{1}{5C_{3}})^{2}{\theta_{\mathrm{min}}^{*}}$. Suppose also that $\delta^{*}_{01}(2s)\leq s\sqrt{nd}/60$, where $\delta^{*}_{01}$ is as in Definition 3.3 with ${F^{*}}$ as in (5.4). Then on an event of probability at least $1-\,\mathrm{exp}(-ds^{2}/C_{1}^{2})$, there exists a unique MLE $\hat{\theta}=\bar{N}\in\Theta$, and it holds
$\begin{split}\chi^{2}(\bar{N}||{\theta^{*}})&\lesssim\frac{s^{2}d}{n},\\ \mathrm{TV}(\gamma_{\ell},\gamma^{*})&\lesssim\frac{sd}{\sqrt{n{\theta_{\mathrm{min}}^{*}}}},\\ \mathrm{TV}(\pi_{\ell},\gamma_{\ell})&\lesssim\frac{d}{\sqrt{n{\theta_{\mathrm{min}}^{*}}}}+e^{-ds^{2}/25C_{1}^{2}},\\ \mathrm{TV}(\pi_{v},\pi_{\ell})&\lesssim\frac{\delta^{*}_{01}(2s)}{\sqrt{n}}+e^{d({M_{0}}^{*}-s^{2}/45C_{1}^{2})}.\end{split}$ (5.9)
Corollary 5.12 (Asymptotic BvM for discrete probability distribution).
Let $d=d_{n}$ and $\Theta_{n}$ be the corresponding sequence of parameter spaces (5.1). Let $\theta^{*}_{n}\in\Theta_{n}$ be a sequence of ground truth pmfs, and write $v_{n},\gamma^{*}_{n},M_{0n}^{*},\delta^{*}_{01n},\theta^{*}_{\mathrm{min},n}$ to emphasize the dependence of these quantities on $n$. If $M_{0n}^{*}=\mathcal{O}(1)$, $\delta^{*}_{01n}(s)=o(\sqrt{n})$ for each $s\geq 0$, and if
$\frac{d_{n}^{2}}{n\theta^{*}_{\mathrm{min},n}}=o(1),$
then $\mathrm{TV}(\pi_{v_{n}},\gamma^{*}_{n})\to 0$ with probability tending to 1 as $n\to\infty$.
This corollary is stronger than the BvM in [6] in two ways. First, we assume only that the prior satisfies $\delta^{*}_{01n}(s)=o(\sqrt{n})$, whereas [6] essentially requires $\delta^{*}_{01n}(s)=o(1)$ provided one defines $\delta^{*}_{01n}(s)$ as a Lipschitz constant; recall the discussion at the end of Section 3.3. Second, we require only that $d_{n}^{2}/n\theta^{*}_{\mathrm{min},n}=o(1)$, whereas [6] requires $d_{n}^{3}/n\theta^{*}_{\mathrm{min},n}=o(1)$.
6 Logistic regression with Gaussian design
In this section, we prove the BvM for logistic regression with Gaussian design. We describe the set-up in Section 4.1 and prove the BvM in Section 4.2. At the end of the section, we compare our result to that of [26] on the BvM for generalized linear models. We note also that similar arguments to the ones below have already appeared in [14] and [13] in the context of Laplace approximation accuracy, but they do not constitute a BvM proof.
6.1 Set-up
In the logistic regression model, we observe feature-label pairs $(X_{i},Y_{i})$, $i=1,\dots,n$, where $X_{i}\in\mathbb{R}^{d}$ is the feature vector and $Y_{i}\in\{0,1\}$ is the label associated to $X_{i}$. We assume the features are drawn as
$X_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,S),\quad i=1,\dots,n,$
where $S\succ 0$ is a nondegenerate covariance matrix. Given a feature $X=x$, the model for the distribution of the corresponding label $Y=y\in\{0,1\}$ is
$p(y\mid\theta,x)=\,\mathrm{exp}\left(yx^{\intercal}\theta-\psi(x^{\intercal}\theta)\right),$ (6.1)
where $\psi(t)=\log(1+e^{t})$. This is just the distribution Bernoulli$(\psi^{\prime}(x^{\intercal}\theta))$, where $\psi^{\prime}(t)=(1+e^{-t})^{-1}$ is the sigmoid function. The parameter $\theta$ is the unknown coefficient vector whose posterior distribution we are interested in. The model (6.1) leads to the following normalized negative log likelihood $\ell$; we also take note of $\ell$’s derivatives.
$\begin{split}\ell(\theta)&=\frac{1}{n}\sum_{i=1}^{n}\left[\psi(X_{i}^{\intercal}\theta)-Y_{i}X_{i}^{\intercal}\theta\right],\\ \nabla\ell(\theta)&=\frac{1}{n}\sum_{i=1}^{n}\left[\psi^{\prime}(X_{i}^{\intercal}\theta)-Y_{i}\right]X_{i},\\ \nabla^{k}\ell(\theta)&=\frac{1}{n}\sum_{i=1}^{n}\psi^{(k)}(X_{i}^{\intercal}\theta)X_{i}^{\otimes k},\quad k\geq 2.\end{split}$ (6.2)
Since $\psi^{\prime\prime}$ is strictly positive, we conclude that $\ell$ is strictly convex as long as the $X_{i}$ span $\mathbb{R}^{d}$. This occurs with probability 1, since we assumed $S\succ 0$. Also, clearly $\ell\in C^{3}(\mathbb{R}^{d})$. Hence Assumption A1 is satisfied. For the weighting matrix, we choose ${F^{*}}=S$. Then ${\mathcal{U}}^{*}(t)=\{\theta\in\mathbb{R}^{d}\;:\;|\theta-{\theta^{*}}|_{S}\leq t\sqrt{d/n}\}$. Also, using equation (6.2) for $\nabla^{3}\ell$, and the fact that $\sup_{t\in\mathbb{R}}|\psi^{\prime\prime\prime}(t)|\leq 1$, we get the following bound on $\delta^{*}_{3}(t)$.
$*.\delta^{*}_{3}(t)=\sup_{\theta\in{\mathcal{U}}^{*}(t)}\sup_{|u|_{S}=1}\left|\frac{1}{n}\sum_{i=1}^{n}\psi^{\prime\prime\prime}(X_{i}^{\intercal}\theta)(X_{i}^{T}u)^{3}\right|\leq\sup_{|u|_{S}=1}\frac{1}{n}\sum_{i=1}^{n}\left|X_{i}^{\intercal}u\right|^{3}=:\delta^{*}_{3}.$ (6.3)
6.2 BvM proof
As for the exponential family model, we prove the BvM by applying Lemma 3.5 together with Lemma 3.7. Recall that the latter lemma shows the event $\bar{E}(s,\lambda)$ defined in (3.7) is a subset of $E(s,\lambda)$ from the former lemma. To finish the BvM proof, we will bound from below the probability of the three events $\{\|\nabla\ell({\theta^{*}})\|_{S}\leq 2\lambda s\sqrt{d/n}\}$, $\{\nabla^{2}\ell({\theta^{*}})\succeq\lambda S\}$, and $\{\delta^{*}_{3}(s)\leq C\}$. The first two are the same as in $\bar{E}(s,\lambda)$, and $\{\delta^{*}_{3}(s)\leq C\}$ implies the third event in $\bar{E}(s,\lambda)$. Furthermore, we can use the bound $\delta^{*}_{3}(s)\leq C$ in the righthand sides of (3.4) and (3.6) to make them deterministic. Note that $\nabla^{2}\ell_{\infty}=\nabla^{2}\ell$. As a result, the term $\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}$ drops out of (3.6) and the condition $\nabla^{2}\ell_{\infty}({\theta^{*}})\succeq\lambda_{\infty}S$ holds on the event $\{\nabla^{2}\ell({\theta^{*}})\succeq\lambda S\}$ with $\lambda=\lambda_{\infty}$.
The key difference between the BvM proof for logistic regression and exponential families is that for logistic regression, $\delta^{*}_{3}$ and $\nabla^{2}\ell_{\infty}$ are both random, and $\nabla^{2}\ell_{\infty}\neq{F^{*}}$. As previously noted, this extra layer of randomness is due to the random design, which we will explicitly account for in our statement of the BvM.
Lemma 6.1 (Adaptation of Lemma 7, Chapter 3, [29]).
Suppose $d<n$. Then for some $\lambda=\lambda(|{\theta^{*}}|_{S})>0$ depending only on $|{\theta^{*}}|_{S}$, the event
$\{\nabla^{2}\ell({\theta^{*}})\succeq\lambda S\}=\left\{\frac{1}{n}\sum_{i=1}^{n}\psi^{\prime\prime}(X_{i}^{\intercal}{\theta^{*}})X_{i}X_{i}^{\intercal}\succeq\lambda S\right\}$
has probability at least $1-4e^{-Cn}$, where $C$ is independent of $\lambda$.
The function $|{\theta^{*}}|_{S}\mapsto\lambda(|{\theta^{*}}|_{S})$ is nonincreasing, so if $|{\theta^{*}}|_{S}$ is bounded above by a constant then $\lambda=\lambda(|{\theta^{*}}|_{S})$ is bounded away from zero by a constant.
Lemma 6.2.
Let $X_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,S)$, $i=1,\dots,n$, and $Y_{i}\mid X_{i}\sim\mathrm{Bernoulli}(\psi^{\prime}(X_{i}^{\intercal}{\theta^{*}}))$. If $d/n\leq 1/8$ and $(\lambda s)^{2}\geq 8$, then the event
$\{\|\nabla\ell({\theta^{*}})\|_{S}\leq 2\lambda s\sqrt{d/n}\}=\left\{\left\|\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\mathbb{E}\,[Y_{i}\mid X_{i}])X_{i}\right\|_{S}\leq 2\lambda s\sqrt{d/n}\right\}$
has probability at least $1-e^{-n/4}-e^{-d(\lambda s)^{2}/4}$.
See Appendix E for the proof. The basic idea is to use an $\epsilon$-net argument, apply Hoeffding’s inequality conditionally on the $X_{i}$, take the expectation over the $X_{i}$ distribution, and use a Gaussian tail bound.
Finally, we bound $\delta^{*}_{3}$ with high probability, recalling from (6.3) that $\delta^{*}_{3}$ is the uniform upper bound on $\delta^{*}_{3}(s)$, $s\geq 0$. In order to bound $\delta^{*}_{3}$, we use the following result.
Theorem 6.3 (Adapted from Theorem 4.2 in [1]).
Let $Z_{1},\dots,Z_{n}$ be i.i.d. standard Gaussian. For every $p\geq 2$ there exist constants $C_{p},C_{p}^{\prime},C_{p}^{\prime\prime}$ depending only on $p$ such that if $C_{p}t^{2p}[\log(2t^{2})]^{2p-2}\leq\left(d^{p/2}/n\right)^{-1}$ then
$\begin{split}\sup_{|u|=1}\frac{1}{n}\sum_{i=1}^{n}|u^{\intercal}Z_{i}|^{p}\leq C_{p}^{\prime}\end{split}$ (6.4)
with probability at least $1-e^{-C_{p}^{\prime\prime}t\sqrt{d}}$.
In a short calculation in Appendix E, we apply the theorem with $p=3$ to show that if $n\geq d^{3/2}$, then
$\mathbb{P}(\delta^{*}_{3}\leq C)\geq 1-e^{-C(nd)^{1/8}}.$ (6.5)
Note that $\{\delta^{*}_{3}\leq C\}$ is a subset of $\{s\delta^{*}_{3}(s)\sqrt{d/n}\leq\lambda/4\}$ provided $s\sqrt{d/n}\leq\lambda/4C$.
By the above results we conclude that $\{\|\nabla\ell({\theta^{*}})\|_{S}\leq 2\lambda s\sqrt{d/n}\}\cap\{\nabla^{2}\ell({\theta^{*}})\succeq\lambda S\}\cap\{\delta^{*}_{3}(s)\leq C\}$ is a subset of $\bar{E}(s,\lambda)$, and has probability at least $1-e^{-C(nd)^{1/8}}-5e^{-Cn}-e^{-(\lambda s)^{2}/4}$, provided
$s\geq\sqrt{8}/\lambda,\quad d/n\leq 1/8,\quad n\geq d^{3/2},\quad s\sqrt{d/n}\leq c\lambda$ (6.6)
for some $c$ small enough. We can now apply Lemma 3.5 to conclude the following BvM.
Corollary 6.4 (Finite-sample BvM for logistic regression).
Let $\lambda=\lambda(|{\theta^{*}}|_{S})$ be as in Lemma 6.1. Let $s,d,n$ satisfy (6.6), and also $s\geq 6\sqrt{2/\lambda}$. Finally, suppose $\delta^{*}_{01}(2s)/\sqrt{nd}\leq\lambda s/48$. Then on an event of probability at least $1-e^{-C(nd)^{1/8}}-5e^{-Cn}-e^{-(\lambda s)^{2}/4}$, there exists a unique MLE $\hat{\theta}$. Furthermore, on this event it holds
$\begin{split}|\hat{\theta}-{\theta^{*}}|_{S}&\lesssim s\sqrt{d/n},\\ \mathrm{TV}(\pi_{v},\pi_{\ell})&\lesssim\lambda^{-1/2}\frac{\delta^{*}_{01}(2s)}{\sqrt{n}}+e^{d[{M_{0}}^{*}-\lambda s^{2}/36]},\\ \mathrm{TV}(\pi_{\ell},\gamma_{\ell})&\lesssim\lambda^{-3/2}\frac{d}{\sqrt{n}}+e^{-d\lambda s^{2}/18},\\ \mathrm{TV}(\gamma_{\ell},\gamma^{*})&\lesssim\lambda^{-1}\frac{sd}{\sqrt{n}}.\end{split}$ (6.7)
We now take $n\to\infty$ and then $s\to\infty$ to prove the following asymptotic BvM.
Corollary 6.5 (Asymptotic BvM for logistic regression).
Let $\theta^{*}_{n}\in\mathbb{R}^{d_{n}}$ be a sequence of ground truth vectors and $S_{n}\succ 0$ be a sequence of positive definite $d_{n}\times d_{n}$ matrices, such that $|\theta^{*}_{n}|_{S_{n}}=\mathcal{O}(1)$ as $n\to\infty$. Write $v_{n},\gamma^{*}_{n},M_{0n}^{*},\delta^{*}_{01n}$ to emphasize the dependence of these quantities on $n$. If $M_{0n}^{*}=\mathcal{O}(1)$, $\delta^{*}_{01n}(s)=o(\sqrt{n})$ for each $s\geq 0$, and
$d_{n}^{2}/n=o(1),$
then $\mathrm{TV}(\pi_{v_{n}},\gamma^{*}_{n})\to 0$ with probability tending to $1$ as $n\to\infty.$ The probability is with respect to the joint feature-label distribution $X_{i}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{N}(0,S_{n})$, $Y_{i}|X_{i}\sim\mathrm{Ber}(\psi^{\prime}(X_{i}^{\intercal}\theta^{*}_{n}))$, $i=1,\dots,n$.
The work [26] proves a BvM for generalized linear models (of which logistic regression is an example), though the design is considered deterministic and the conditions on the design under which the BvM holds are left unspecified. For a comparison of the prior conditions, see Section 3.3. Finally, the key difference is that [26] requires $d_{n}^{3}/n=o(1)$ for the BvM to hold.
7 Discussion
We have presented the proof of the BvM for models of varying complexity, thereby demonstrating that the same basic template can be used for a range of models. In each of the three models, our proof of the BvM with rate $d^{2}/n$ centers around the three bounds in Lemma 3.5. The improved dimension dependence stems from the middle bound on $\mathrm{TV}(\pi_{\ell},\gamma_{\ell})$, which is the Laplace approximation error and is controlled by our main Theorem 2.2. We anticipate that the same proof structure can be used to prove the BvM with the improved rate for more general statistical models with either log concave likelihoods or log concave posteriors. In the case of log concave posteriors (not considered in the present work), it is likely possible to chart a similar course: to bound $\mathrm{TV}(\pi_{v},\gamma_{v})$ using Theorem 2.2, and bound $\mathrm{TV}(\gamma_{v},\gamma^{*})$ by controlling the distance between the MAP and MLE and the distance between the Hessian of $v$ at the MAP and the Fisher information matrix. We leave this case to future work.
Although BvM results from prior works have suboptimal dimension dependence, certainly many of the proof techniques developed in these works can be incorporated into a proof with the optimal rate. By no means do we claim that our template based on the Laplace approximation is the only way to achieve the $d^{2}/n$ rate.
Appendix A Proofs from Section 2
Recall that in this section, $\delta_{3}(r)$ is defined as
$\delta_{3}(r)=\sup_{|\theta-\hat{\theta}|_{H}\leq r\sqrt{d/n}}\frac{\|\nabla^{2}f(\theta)-\nabla^{2}f(\hat{\theta})\|_{H}}{|\theta-\hat{\theta}|_{H}},\quad H=\nabla^{2}f(\hat{\theta}).$ (A.1)
A.1 Proofs from Section 2.1
Proof of Lemma 2.8.
To prove the lower bound on $\nabla^{2}f(\theta)$ in ${\mathcal{U}}(r)$, note that
$\|\nabla^{2}f(\theta)-H\|_{H}=\|\nabla^{2}f(\theta)-\nabla^{2}f(\hat{\theta})\|_{H}\leq\delta_{3}(r)\|\theta-\hat{\theta}\|_{H}=r\sqrt{d/n}\delta_{3}(r)\leq 1/2,$ (A.2)
using the assumption on $\delta_{3}(r)$ from Theorem 2.2. Hence $\nabla^{2}f(\theta)\succeq\frac{1}{2}H$ for all $\theta\in{\mathcal{U}}(r)$. To prove the linear growth bound, recall that $f$ is convex by assumption. Fix a point $\theta\in\Theta$ such that $|\theta-\hat{\theta}|_{H}\geq r\sqrt{d/n}$. Then the whole segment between $\hat{\theta}$ and $\theta$ is also in $\Theta$. In particular, convexity of $f$ gives
$\begin{split}f(\theta)-f(\hat{\theta})\geq\frac{1}{t}\left[f(\hat{\theta}+t(\theta-\hat{\theta}))-f(\hat{\theta})\right],\quad t=\frac{r\sqrt{d/n}}{|\theta-\hat{\theta}|_{H}}.\end{split}$
Dividing both sides by $|\theta-\hat{\theta}|_{H}$ , we get
$\begin{split}\frac{f(\theta)-f(\hat{\theta})}{|\theta-\hat{\theta}|_{H}}\geq\frac{f(\hat{\theta}+t(\theta-\hat{\theta}))-f(\hat{\theta})}{|t(\theta-\hat{\theta})|_{H}}\geq\inf_{|u|_{H}=r\sqrt{d/n}}\frac{f(\hat{\theta}+u)-f(\hat{\theta})}{|u|_{H}}=:\psi.\end{split}$ (A.3)
To get the second inequality, we replaced $t(\theta-\hat{\theta})$, which has $H$-norm $r\sqrt{d/n}$ by construction, with any $u$ such that $|u|_{H}=r\sqrt{d/n}$. To finish the proof, we bound $\psi$ from below. A Taylor expansion around $\theta=\hat{\theta}$ gives that for $|u|_{H}=r\sqrt{d/n}$, we have
$\begin{split}f(\hat{\theta}+u)-f(\hat{\theta})&=\frac{1}{2}\langle\nabla^{2}f(\hat{\theta}+\xi),u^{\otimes 2}\rangle\geq\frac{1}{4}|u|_{H}^{2},\end{split}$ (A.4)
using that $\nabla^{2}f(\theta)\succeq\frac{1}{2}H$ for all $\theta\in{\mathcal{U}}(r)$ to get the second inequality. We now divide by $|u|_{H}$ to get
$\begin{split}\frac{f(\theta)-f(\hat{\theta})}{|u|_{H}}&\geq\frac{1}{4}|u|_{H}=\frac{r}{4}\sqrt{d/n}.\end{split}$ (A.5)
Substituting this bound into (A.3) concludes the proof. ∎
Proof of Lemma 2.9.
We let ${\mathcal{U}}^{c}$ be the complement of ${\mathcal{U}}$ in $\mathbb{R}^{d}$, and omit the range of integration when the integral is over all of $\mathbb{R}^{d}$. First, we have
$\begin{split}2\mathrm{TV}(\mu,\hat{\mu})=\int\left|\frac{F}{\int F}-\frac{\hat{F}}{\int\hat{F}}\right|\leq\frac{\int_{{\mathcal{U}}^{c}}F}{\int F}+\frac{\int_{{\mathcal{U}}^{c}}\hat{F}}{\int\hat{F}}+\int_{\mathcal{U}}\left|\frac{F}{\int F}-\frac{\hat{F}}{\int\hat{F}}\right|\end{split}$ (A.6)
Next, note that
$\begin{split}\int_{\mathcal{U}}&\left|\frac{F}{\int F}-\frac{\hat{F}}{\int\hat{F}}\right|-\int_{\mathcal{U}}\left|\frac{F}{\int_{\mathcal{U}}F}-\frac{\hat{F}}{\int_{\mathcal{U}}\hat{F}}\right|\leq\int_{\mathcal{U}}\left|\frac{F}{\int F}-\frac{F}{\int_{\mathcal{U}}F}\right|+\int_{\mathcal{U}}\left|\frac{\hat{F}}{\int_{\mathcal{U}}\hat{F}}-\frac{\hat{F}}{\int\hat{F}}\right|\\ &=\int_{\mathcal{U}}F\left|\left(\int_{\mathcal{U}}F\right)^{-1}-\left(\int F\right)^{-1}\right|+\int_{\mathcal{U}}\hat{F}\left|\left(\int_{\mathcal{U}}\hat{F}\right)^{-1}-\left(\int\hat{F}\right)^{-1}\right|\\ &=\left|1-\frac{\int_{\mathcal{U}}F}{\int F}\right|+\left|1-\frac{\int_{\mathcal{U}}\hat{F}}{\int\hat{F}}\right|=\frac{\int_{{\mathcal{U}}^{c}}F}{\int F}+\frac{\int_{{\mathcal{U}}^{c}}\hat{F}}{\int\hat{F}}\end{split}$ (A.7)
Substituting (A.7) into (A.6) gives
$\begin{split}2\mathrm{TV}(\mu,\hat{\mu})&\leq 2\frac{\int_{{\mathcal{U}}^{c}}F}{\int F}+2\frac{\int_{{\mathcal{U}}^{c}}\hat{F}}{\int\hat{F}}+\int_{\mathcal{U}}\left|\frac{F}{\int_{\mathcal{U}}F}-\frac{\hat{F}}{\int_{\mathcal{U}}\hat{F}}\right|\\ &=2\frac{\int_{{\mathcal{U}}^{c}}F}{\int_{\mathcal{U}}F}+2\frac{\int_{{\mathcal{U}}^{c}}\hat{F}}{\int_{\mathcal{U}}\hat{F}}+2\mathrm{TV}\left(\mu|_{\mathcal{U}},\hat{\mu}|_{\mathcal{U}}\right).\end{split}$ (A.8)
Dividing by 2 finishes the proof of the first statement.
Next we prove the second statement. For brevity, redefine $\mu$ and $\hat{\mu}$ to be their restrictions to ${\mathcal{U}}$ and recall that $\mu\propto e^{-nf}$, $\hat{\mu}\propto e^{-n\hat{f}}$ on ${\mathcal{U}}$. Let $T(x)=(nH)^{1/2}x$. Then
$(T_{\#}\mu)(y)\propto e^{-f_{H}(y)},\quad(T_{\#}\hat{\mu})(y)\propto e^{-\hat{f}_{H}(y)},\quad y\in(nH)^{1/2}{\mathcal{U}},$
where
$f_{H}(y)=nf((nH)^{-1/2}y),\quad\hat{f}_{H}(y)=n\hat{f}((nH)^{-1/2}).$
Note that $\nabla^{2}f_{H}(y)=H^{-1/2}\nabla^{2}f((nH)^{-1/2}y)H^{-1/2}\succeq\lambda I_{d}$ for all $y\in(nH)^{1/2}{\mathcal{U}}$ by the assumption $\nabla^{2}f(x)\succeq\lambda H$ for all $x\in{\mathcal{U}}$. Therefore, $T_{\#}\mu$ is $\lambda$-strongly concave, so it satisfies a log-Sobolev inequality (LSI) with constant $1/\lambda$ [2]. Using the affine invariance of TV distance, then Pinsker’s inequality, then the LSI, we get
$\mathrm{TV}(\hat{\mu},\mu)^{2}=\mathrm{TV}(T_{\#}\hat{\mu},T_{\#}\mu)^{2}\leq\frac{1}{2}\mathrm{KL}\left(\,T_{\#}\hat{\mu}\;||\;T_{\#}\mu\right)\leq\frac{1}{4\lambda}\mathrm{FI}\left(\,T_{\#}\hat{\mu}\;||\;T_{\#}\mu\right)$ (A.9)
where
$\mathrm{FI}\left(\,T_{\#}\hat{\mu}\;||\;T_{\#}\mu\right)=\mathbb{E}\,_{Y\sim T_{\#}\hat{\mu}}\left[\left|\nabla\log\frac{T_{\#}\hat{\mu}}{T_{\#}\mu}(Y)\right|^{2}\right]$
Now, $\log((T_{\#}\hat{\mu}/T_{\#}\mu)(Y))=(f_{H}-\hat{f}_{H})(Y)=n(f-\hat{f})((nH)^{-1/2}Y).$ Therefore,
$\begin{split}\nabla\log((T_{\#}\hat{\mu}/T_{\#}\mu)(Y))&=\sqrt{n}H^{-1/2}\left(\nabla(f-\hat{f})((nH)^{-1/2}Y)\right)\\ &\stackrel{{\scriptstyle d}}{{=}}\sqrt{n}H^{-1/2}\nabla(f-\hat{f})(X),\end{split}$
where $X\sim\hat{\mu}$. Therefore,
$\mathrm{FI}\left(\,T_{\#}\hat{\mu}\;||\;T_{\#}\mu\right)=n\mathbb{E}\,_{X\sim\hat{\mu}}\left[\left\|\nabla(f-\hat{f})(X)\right\|_{H}^{2}\right].$
Substituting this into (A.9) finishes the proof. ∎
A.2 Proof of Theorem 2.2
In this section, we make the assumptions from Theorem 2.2. Recall that it suffices to bound the local expectation and the two tail integrals in (2.10). These quantities are bounded in Lemma A.1 and Corollary A.5 below. The proof of Proposition 2.2 follows immediately by adding up these bounds.
A.2.1 Local Expectation
Lemma A.1.
It holds
$\sqrt{\frac{n}{2}}\mathbb{E}\,_{\theta\sim\gamma_{f}|{\mathcal{U}}}\left[\|\nabla f(\theta)-H(\theta-\hat{\theta})\|_{H}^{2}\right]^{\frac{1}{2}}\leq\frac{\delta_{3}(r)}{\sqrt{2}}\frac{d}{\sqrt{n}}.$ (A.10)
Proof.
First note
[…]
Hence if $\theta\in{\mathcal{U}}(r)$ then
[…]
Therefore, using that $\hat{\theta}+(nH)^{-1/2}Z\sim\gamma_{f}$ when $Z\sim\mathcal{N}(0,I_{d})$, we get
$\begin{split}\mathbb{E}\,_{\theta\sim\gamma_{f}|{\mathcal{U}}}&\left[\|\nabla f(\theta)-H(\theta-\hat{\theta})\|_{H}^{2}\right]\leq\frac{1}{4}\delta_{3}(r)^{2}\mathbb{E}\,_{\theta\sim\gamma_{f}|{\mathcal{U}}}[\|\theta-\hat{\theta}\|_{H}^{4}]\\ &=\frac{1}{4n^{2}}\mathbb{P}(|Z|\leq r\sqrt{d})^{-1}\mathbb{E}\,\left[|Z|^{4}\mathds{1}(|Z|\leq r\sqrt{d})\right]\leq\frac{3\delta_{3}(r)^{2}d^{2}}{4n^{2}}\mathbb{P}(|Z|\leq r\sqrt{d})^{-1}\end{split}$ (A.11)
Now note that
$\mathbb{P}(|Z|\leq r\sqrt{d})\geq 1-e^{-d(r-1)^{2}/2}\geq 1-e^{-25/2}\geq 3/4$
for all $d\geq 1$ since $r\geq 6$. Substituting this bound into (A.11) gives
$\mathbb{E}\,_{\theta\sim\gamma_{f}|{\mathcal{U}}}\left[\|\nabla f(\theta)-H(\theta-\hat{\theta})\|_{H}^{2}\right]\leq\delta_{3}(r)^{2}\frac{d^{2}}{n^{2}}$ (A.12)
Taking the square root and then multiplying by $\sqrt{n/2}$ gives the desired bound. ∎
A.2.2 Tail integrals
Lemma A.2.
Let ${\mathcal{U}}={\mathcal{U}}(r)$ for some $r>2$. It holds
[…]
Proof.
(2.6) gives
[…]
Using Lemma F.3 with $a=r$ and $b=r/4$ (the lemma applies since we assumed $r^{2}/4>1$), we get
[…]
To conclude, bound $de^{3d/2}$ by $e^{2d}$, noting that $de^{-d/2}\leq 1$ for all $d$. ∎
Lemma A.3.
Let ${\mathcal{U}}={\mathcal{U}}(r)$ for some $r\geq 2$. Then
[…]
Proof.
For all $|\theta-\hat{\theta}|_{H}\leq r\sqrt{d/n}$, it holds
$\begin{split}f(\theta)-f(\hat{\theta})&=\frac{1}{2}\left\langle\nabla^{2}f(\xi),(\theta-\hat{\theta})^{\otimes 2}\right\rangle\leq\frac{3}{4}|\theta-\hat{\theta}|_{H}^{2},\end{split}$ (A.17)
using that $\sup_{\theta\in{\mathcal{U}}}\|\nabla^{2}f(\theta)\|_{H}\leq 3/2$, since $\|\nabla^{2}f(\theta)-H\|_{H}\leq 1/2$, shown in the proof of Lemma 2.8. Hence
[…]
To conclude the proof note that when $r\geq 2$ we have $\mathbb{P}(|Z|\leq r\sqrt{3/2}\sqrt{d})\geq\mathbb{P}\left(|Z|\leq\sqrt{6}\sqrt{d}\right)\geq 1-e^{-d(\sqrt{6}-1)^{2}/2}\geq 1/2$ for all $d\geq 1$. ∎
Corollary A.4.
Let ${\mathcal{U}}={\mathcal{U}}(r)$. If $r\geq 6$ then
[…]
Proof.
Dividing the upper bound on the numerator by the lower bound on the denominator gives
[…]
We now note that $2+\log(\sqrt{3/2})\leq 2.21$, and that
$2.21+\log r\leq\frac{r^{2}}{9}\quad\forall r\geq 6$ (A.19)
Substituting this bound into the exponential concludes the proof. ∎
Corollary A.5.
Let $r\geq 6$. Then
$T:=\frac{\int_{\Theta\setminus{\mathcal{U}}}e^{-nf}}{\int_{\mathcal{U}}e^{-nf}}+\gamma_{f}({\mathcal{U}}^{c})\leq 3\,\mathrm{exp}\left(-dr^{2}/9\right).$
Proof.
From the above corollary and a standard Gaussian tail bound (i.e. using that $\gamma_{f}({\mathcal{U}}(r)^{c})=\mathbb{P}(|Z|\geq r\sqrt{d})\leq\,\mathrm{exp}(-d(r-1)^{2}/2)$) we get
$\begin{split}T\leq 2e^{-dr^{2}/9}+\,\mathrm{exp}\left(-\frac{d}{2}\left(r-1\right)^{2}\right)\leq 3e^{-dr^{2}/9},\end{split}$ (A.20)
since $(d/2)(r-1)^{2}\geq(d/2)(r/2)^{2}=dr^{2}/8$. ∎
A.3 Analysis of the bound from [13]
We begin by describing the rescaling which turns [13]’s bound (2.11) into the affine-invariant bound (2.13). Define $\bar{f}(\bar{\theta})=(f\circ T^{-1})(\bar{\theta})$, where $T^{-1}(\bar{\theta})=\hat{\theta}+H_{f}^{-1/2}\bar{\theta}$. As discussed in Remark 2.7, it holds $\mathrm{TV}(\pi_{f},\gamma_{f})=\mathrm{TV}(\pi_{\bar{f}},\gamma_{\bar{f}})$. Therefore, we can apply the bound (2.11) to $\bar{f}$. The new domain is $\bar{\Theta}=H_{f}^{1/2}(\Theta-\hat{\theta})$, and $\bar{f}$ is minimized at 0, with $H_{\bar{f}}=I_{d}$. We have
$M_{\bar{f}}=\sup_{|\bar{\theta}|\leq r\sqrt{d/n}}\|\nabla^{3}\bar{f}(\bar{\theta})\|=\sup_{|\theta-\hat{\theta}|_{H}\leq r\sqrt{d/n}}\|\nabla^{3}f(\theta)\|_{H}=\delta_{3}(r),$
and hence $H_{\bar{f}}^{p}=\left(1+\frac{r\delta_{3}(r)}{3}\sqrt{d/n}\right)I_{d}$. Next,
$\kappa_{\bar{f}}=\inf_{|\bar{\theta}|\geq r\sqrt{d/n}}\bar{f}(\bar{\theta})-\bar{f}(0)=\inf_{|\bar{\theta}|\geq r\sqrt{d/n}}f(\hat{\theta}+H_{f}^{-1/2}\bar{\theta})-f(\hat{\theta})=\inf_{|\theta-\hat{\theta}|_{H_{f}}\geq r\sqrt{d/n}}f(\theta)-f(\hat{\theta}),$
which was denoted simply $\kappa$ in (2.14). Finally, we have $\mathscr{D}_{\bar{f}}=\,\mathrm{exp}(-\frac{1}{2}(r\sqrt{d}-\sqrt{d})^{2})=\,\mathrm{exp}(-\frac{d}{2}(r-1)^{2})$ and similarly
$\mathscr{D}_{\bar{f}}^{p}=\,\mathrm{exp}\left(-\frac{1}{2}\left(r\sqrt{d}-\sqrt{d/\lambda}\right)^{2}\lambda\right)=\,\mathrm{exp}\left(-\frac{d}{2}(r\sqrt{\lambda}-1)^{2}\right),\qquad\lambda=1+\frac{r\delta_{3}(r)}{3}\sqrt{d/n}.$
Using all of these expressions in (2.11) gives
$\begin{split}\mathrm{TV}(\pi_{f},\gamma_{f})=\mathrm{TV}(\pi_{\bar{f}},\gamma_{\bar{f}})\leq&\frac{\delta_{3}(r)d}{4\sqrt{(1-r\delta_{3}(r)\sqrt{d/n})(1-e^{-d(r-1)^{2}/2})}\sqrt{n}}+2e^{-\frac{d}{2}(r-1)^{2}}\\ &+2\frac{\left(1+\frac{r\delta_{3}(r)}{3}\sqrt{d/n}\right)^{d/2}\mathrm{Vol}(H_{f}^{1/2}\Theta)}{(2\pi)^{d/2}(1-e^{-d(r\sqrt{\lambda}-1)^{2}/2})}n^{d/2}e^{-n\kappa}\end{split}$ (A.21)
We assume $r$ is large enough that $01-e^{-d(r-1)^{2}/2}\geq c>$ 0 for some absolute constant $c$; otherwise the bound blows up. Note that $\sqrt{\lambda}\geq 1$, hence $1-e^{-d(r\sqrt{\lambda}-1)^{2}/2}>c$ as well. Similarly, we assume $r\delta_{3}(r)\sqrt{d/n}\leq 1-c$. Using this bound everywhere $r\delta_{3}(r)\sqrt{d/n}$ appears, (A.21) simplifies to (2.13).
Finally, we prove the upper bound on $\kappa$. First note that $\|\nabla^{2}f(\theta)\|_{H_{f}}\leq 2$ for all $\theta\in{\mathcal{U}}(r)$ using Taylor’s theorem, the fact that $\|\nabla^{2}f(\hat{\theta})\|_{H_{f}}=1$, and the bound $r\delta_{3}(r)\sqrt{d/n}<1$.
Next, fix $\theta$ on the boundary of ${\mathcal{U}}(r)$, i.e. $|\theta-\hat{\theta}|_{H_{f}}=r\sqrt{d/n}$. Then again by Taylor’s theorem and the fact that $\|\nabla^{2}f\|_{H_{f}}\leq 2$ on ${\mathcal{U}}(r)$, we have
$f(\theta)-f(\hat{\theta})=\frac{1}{2}(\theta-\hat{\theta})^{\intercal}\nabla^{2}f(\xi)(\theta-\hat{\theta})\leq\frac{1}{2}|\theta-\hat{\theta}|_{H_{f}}^{2}\|\nabla^{2}f(\xi)\|_{H_{f}}\leq(r\sqrt{d/n})^{2}=r^{2}d/n.$
Thus $\kappa\leq\inf_{|\theta-\hat{\theta}|_{H_{f}}=r\sqrt{d/n}}f(\theta)-f(\hat{\theta})\leq r^{2}d/n$.
Appendix B Proofs from Section 3.3
Definition B.1.
On the event $\{\exists!\,\mathrm{MLE}\,\hat{\theta}\}$, define $H=\nabla^{2}\ell(\hat{\theta})$ and ${\mathcal{U}}(r)=\{|\theta-\hat{\theta}|_{H}\leq r\sqrt{d/n}\}$. Assuming ${\mathcal{U}}(r)\subset\Theta$, define also $\delta_{3}(r)=\sup_{\theta\in{\mathcal{U}}(r)}\|\nabla^{3}\ell(\theta)\|_{H}$.
Lemma B.2.
On the event $E(s,\lambda)=\{\exists!\,\mathrm{MLE}\,\hat{\theta}\}\cap\{\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}\}\cap\{|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}\}\cap\{2s\delta^{*}_{3}(2s)\sqrt{d/n}\leq\lambda/2\}$, it holds
$\nabla^{2}\ell(\hat{\theta})\succeq\tau{F^{*}},\quad\tau=\lambda/2,$
$\begin{gathered}{\mathcal{U}}(\sqrt{\tau}s)\subseteq{\mathcal{U}}^{*}(2s)\qquad\delta_{3}(\sqrt{\tau}s)\leq\tau^{-3/2}\delta^{*}_{3}(2s),\end{gathered}$ (B.1)
$\sqrt{\tau}s\delta_{3}(\sqrt{\tau}s)\sqrt{d/n}\leq 1/2.$ (B.2)
See Section B.1 for the proof.
Proof of Bound (3.4).
By Assumption A1, $\ell$ is convex and $\ell\in C^{3}(\Theta)$ with probability 1. Also, $\ell$ has a unique global minimizer $\hat{\theta}$ on $E(s,\lambda)$. Let $r=\sqrt{\tau}s$ where $\tau=\lambda/2$, so that $r\geq 6$. By Lemma B.2, it holds on $E(s,\lambda)$ that ${\mathcal{U}}(r)\subset{\mathcal{U}}^{*}(2s)\subset\Theta$ and $r\delta_{3}(r)\sqrt{d/n}\leq 1/2$. Therefore, we can apply Theorem 2.2 with $f=\ell$ to get that $\mathrm{TV}(\pi_{\ell},\gamma_{\ell})\leq\delta_{3}(r)\frac{d}{\sqrt{2n}}+3e^{-dr^{2}/9}$. Using that $\delta_{3}(r)\leq\tau^{-3/2}\delta^{*}_{3}(2s)$ on $E(s,\tau)$ and that $e^{-dr^{2}/9}=e^{-d\tau s^{2}/9}$ concludes the proof.∎
Proof of Bound (3.5).
Let $\tau=\lambda/2$ and $r=\sqrt{\tau}s$, so that $r\geq 6$. Since $22s\delta^{*}_{3}(2s)\sqrt{d/n}\leq\lambda/$ 2 and $\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}$ on $E(s,\lambda)$, it follows that $\nabla^{2}\ell(\theta)\succeq\frac{\lambda}{2}{F^{*}}$ for all $\theta\in{\mathcal{U}}^{*}(2s)$. We apply Lemma 2.9 with ${\mathcal{U}}={\mathcal{U}}^{*}(2s)$, $\mu=\pi_{\ell}$, $\hat{\mu}=\pi_{v}$, $\lambda=\lambda/2$, and ${F^{*}}$ instead of $H$. This gives
$\mathrm{TV}(\pi_{v},\pi_{\ell})\leq\frac{\int_{\Theta\setminus{\mathcal{U}}^{*}(2s)}e^{-n\ell}\pi_{0}}{\int_{{\mathcal{U}}^{*}(2s)}e^{-n\ell}\pi_{0}}+\frac{\int_{\Theta\setminus{\mathcal{U}}^{*}(2s)}e^{-n\ell}}{\int_{{\mathcal{U}}^{*}(2s)}e^{-n\ell}}+\frac{\sqrt{n}}{\sqrt{2\lambda}}\mathbb{E}\,_{\theta\sim\pi_{v}|_{\mathcal{U}}}\left[\|\nabla(n^{-1}v_{0})\|_{F^{*}}^{2}\right]^{\frac{1}{2}}.$ (B.3)
In the last term, we simply bound the square root of the expectation of the square by the maximum of $\|\nabla(n^{-1}v_{0}(\theta))\|_{F^{*}}$ over $\theta\in{\mathcal{U}}^{*}(2s)$, which is $n^{-1}\delta^{*}_{01}(2s)$ by definition. Thus
$\frac{\sqrt{n}}{\sqrt{2\lambda}}\mathbb{E}\,_{\theta\sim\nu|_{\mathcal{U}}}\left[\|\nabla(n^{-1}v_{0})\|_{F^{*}}^{2}\right]^{\frac{1}{2}}\leq\frac{1}{\sqrt{2n\lambda}}\delta^{*}_{01}(2s).$ (B.4)
Now, by definition of ${M_{0}}^{*}$ and $\delta^{*}_{01}$ and using that ${\mathcal{U}}(r)\subset{\mathcal{U}}^{*}(2s)$ on $E(s,\lambda)$ by Lemma B.2, we have
$\begin{split}\frac{\int_{{\mathcal{U}}^{*}(2s)^{c}}e^{-n\ell}\pi_{0}}{\int_{{\mathcal{U}}^{*}(2s)}e^{-n\ell}\pi_{0}}&\leq\frac{\sup_{\theta\in\Theta}\pi_{0}(\theta)/\pi_{0}({\theta^{*}})}{\inf_{\theta\in{\mathcal{U}}^{*}(2s)}\pi_{0}(\theta)/\pi_{0}({\theta^{*}})}\frac{\int_{{\mathcal{U}}^{*}(2s)^{c}}e^{-n\ell}}{\int_{{\mathcal{U}}^{*}(2s)}e^{-n\ell}}\\ &\leq\,\mathrm{exp}(d{M_{0}}^{*}+2s\delta^{*}_{01}(2s)\sqrt{d/n})\frac{\int_{{\mathcal{U}}^{*}(2s)^{c}}e^{-n\ell}}{\int_{{\mathcal{U}}^{*}(2s)}e^{-n\ell}}\\ &\leq\,\mathrm{exp}(d{M_{0}}^{*}+2s\delta^{*}_{01}(2s)\sqrt{d/n})\frac{\int_{{\mathcal{U}}(r)^{c}}e^{-n\ell}}{\int_{{\mathcal{U}}(r)}e^{-n\ell}}\end{split}$ (B.5)
Hence
$\begin{split}\frac{\int_{\Theta\setminus{\mathcal{U}}^{*}(2s)}e^{-n\ell}\pi_{0}}{\int_{{\mathcal{U}}^{*}(2s)}e^{-n\ell}\pi_{0}}&+\frac{\int_{\Theta\setminus{\mathcal{U}}^{*}(2s)}e^{-n\ell}}{\int_{{\mathcal{U}}^{*}(2s)}e^{-n\ell}}\leq 2\,\mathrm{exp}\left(d[{M_{0}}^{*}+2s\delta^{*}_{01}(2s)/\sqrt{nd}]\right)\frac{\int_{\Theta\setminus{\mathcal{U}}(r)}e^{-n\ell}}{\int_{{\mathcal{U}}(r)}e^{-n\ell}}\\ \leq 4\,\mathrm{exp}\bigg{(}d&\left[{M_{0}}^{*}+2s\delta^{*}_{01}(2s)/\sqrt{nd}-5r^{2}/36\right]\bigg{)}\leq 4\,\mathrm{exp}\left(d({M_{0}}^{*}-s^{2}\lambda/36)\right).\end{split}$ (B.6)
We used (B.5) to get the first inequality, Corollary A.4 to get the second inequality, and that $242s\delta^{*}_{01}(2s)/\sqrt{nd}\leq s^{2}\lambda/$ 24, $725r^{2}/36=5s^{2}\lambda/$ 72 to get the third inequality. Combining (B.6) with (B.4) concludes the proof.
Proof of Bound (3.6).
We use Lemma F.2 with $A_{1}=\nabla^{2}\ell(\hat{\theta})$, $A_{2}=\nabla^{2}\ell_{\infty}({\theta^{*}})$, $\tau_{1}=\lambda$, and $\tau_{2}=\lambda_{\infty}$. Note that
$\|A_{1}-A_{2}\|_{F^{*}}\leq\|\nabla^{2}\ell(\hat{\theta})-\nabla^{2}\ell({\theta^{*}})\|_{F^{*}}+\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}\leq s\delta^{*}_{3}(s)\sqrt{d/n}+\|\nabla^{2}(\ell-\ell_{\infty})({\theta^{*}})\|_{F^{*}}=:\epsilon$
on $E(s,\tau)$. Substituting the above values of $\tau_{1},\tau_{2},\epsilon$ into (F.3) concludes the proof of (3.6). ∎
Remark B.3.
The above proof shows that (3.6) holds with $\delta^{*}_{3}(s)$ in place of $\delta^{*}_{3}(2s)$. In the statement of the lemma, we present the bound using $\delta^{*}_{3}(2s)$ because this slightly larger quantity is what appears in (3.5).
Proof of Lemma 3.7.
We check the assumptions of Corollary F.6 with $h=\ell$, $y_{0}={\theta^{*}}$, and $q=s\sqrt{d/n}$. The condition $\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}$ and $\|\nabla\ell({\theta^{*}})\|_{F^{*}}\leq 2\lambda q$ are both satisfied on the event in the statement of Lemma 3.7. The condition $\|\nabla^{2}\ell(\theta)-\nabla^{2}\ell({\theta^{*}})\|_{F^{*}}\leq\lambda/4$ is also satisfied on this event, since
$\begin{split}\|\nabla^{2}\ell(\theta)-\nabla^{2}\ell({\theta^{*}})\|_{F^{*}}&\leq s\delta^{*}_{3}(s)\sqrt{d/n}\leq\lambda/4.\end{split}$ (B.7)
Therefore, the assumptions of Corollary F.6 are satisfied so we conclude there exists $\hat{\theta}$ such that $\nabla\ell(\hat{\theta})=0$ and $|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq q=s\sqrt{d/n}$. Also, we immediately get by (B.7) that $\|\nabla^{2}\ell(\hat{\theta})-\nabla^{2}\ell({\theta^{*}})\|_{F^{*}}\leq\lambda/4$. Using this and the fact that $\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}$ we get $\nabla^{2}\ell(\hat{\theta})\succeq(3\lambda/4){F^{*}}$. Hence $\hat{\theta}$ is a strict local minimizer, and therefore a global minimizer, since $\ell$ is convex.∎
Proof of bounds in Example 3.8.
The bounds for the flat prior are trivial. For the Gaussian prior $\pi_{0}=\mathcal{N}(\mu,\Sigma)$, we have $\log\pi_{0}(\theta)=-\frac{1}{2}(\theta-\mu)^{\intercal}\Sigma^{-1}(\theta-\mu)+\mathrm{const}$, so that
$\begin{split}d{M_{0}}^{*}=\sup_{\theta\in\mathbb{R}^{d}}\log\pi_{0}(\theta)-\log\pi_{0}({\theta^{*}})=\frac{1}{2}({\theta^{*}}-\mu)^{\intercal}\Sigma^{-1}({\theta^{*}}-\mu)=\frac{1}{2}|{\theta^{*}}-\mu|_{\Sigma^{-1}}^{2}\end{split}$ (B.8)
Next, we have $\nabla\log\pi_{0}(\theta)=\Sigma^{-1}(\theta-\mu)$, and therefore
$\|\nabla\log\pi_{0}(\theta)\|_{F^{*}}=|{F^{*}}^{-1/2}\Sigma^{-1}(\theta-\mu)|\leq\|\Sigma^{-1}\|_{F^{*}}|\theta-\mu|_{F^{*}}.$
Hence
$\delta^{*}_{01}(r)\leq\|\Sigma^{-1}\|_{F^{*}}(|\mu-{\theta^{*}}|_{F^{*}}+r\sqrt{d/n}).$
For the multivariate student’s t prior $\pi_{0}=t_{\nu}(\mu,\Sigma)$, we have $\log\pi_{0}(\theta)=-\frac{\nu+d}{2}\log\left(1+\frac{1}{\nu}|\theta-\mu|^{2}_{\Sigma^{-1}}\right)+\mathrm{const}$, so
$\begin{split}\log\pi_{0}(\theta)&-\log\pi_{0}({\theta^{*}})=\frac{\nu+d}{2}\log\left(\frac{1+\frac{1}{\nu}|{\theta^{*}}-\mu|^{2}_{\Sigma^{-1}}}{1+\frac{1}{\nu}|\theta-\mu|^{2}_{\Sigma^{-1}}}\right)\\ &\leq\frac{\nu+d}{2}\log\left(1+\frac{1}{\nu}|{\theta^{*}}-\mu|^{2}_{\Sigma^{-1}}\right)\leq\frac{\nu+d}{2\nu}|{\theta^{*}}-\mu|^{2}_{\Sigma^{-1}}.\end{split}$ (B.9)
Hence ${M_{0}}^{*}\leq\frac{\nu+d}{2\nu d}|{\theta^{*}}-\mu|^{2}_{\Sigma^{-1}}$. Finally,
$\nabla\log\pi_{0}(\theta)=-\frac{\nu+d}{\nu}\frac{\Sigma^{-1}(\theta-\mu)}{1+\frac{1}{\nu}|\theta-\mu|^{2}_{\Sigma^{-1}}}$
so
$\begin{split}\delta^{*}_{01}(r)&=\sup_{\theta\in{\mathcal{U}}(r)}\|\nabla\log\pi_{0}(\theta)\|_{F^{*}}\leq\frac{\nu+d}{\nu}\sup_{\theta\in{\mathcal{U}}(r)}\|\Sigma^{-1}(\theta-\mu)\|_{{F^{*}}}\\ &\leq\frac{\nu+d}{\nu}\|\Sigma^{-1}\|_{F^{*}}(|\mu-{\theta^{*}}|_{F^{*}}+r\sqrt{d/n})\end{split}$ (B.10)
B.1 Auxiliary results
Proof of Lemma B.2.
Using the definition of event $E(s,\lambda)$ and the definition of $\delta^{*}_{3}$, we have
$\|\nabla^{2}\ell(\hat{\theta})-\nabla^{2}\ell({\theta^{*}})\|_{F^{*}}\leq s\delta^{*}_{3}(s)\sqrt{d/n}\leq\lambda/2.$
Since $\nabla^{2}\ell({\theta^{*}})\succeq\lambda{F^{*}}$, we conclude that $\nabla^{2}\ell(\hat{\theta})\succeq\tau{F^{*}}$ for $\tau=\lambda/2$. This proves the first statement. Next, fix $\theta\in{\mathcal{U}}(\sqrt{\tau}s)$, so that $|\theta-\hat{\theta}|_{H}\leq\sqrt{\tau}s\sqrt{d/n}$. Using that $H=\nabla^{2}\ell(\hat{\theta})\succeq\tau{F^{*}}$ and $|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq s\sqrt{d/n}$ on $E(s,\lambda)$, we then have
$|\theta-{\theta^{*}}|_{F^{*}}\leq|\theta-\hat{\theta}|_{F^{*}}+|\hat{\theta}-{\theta^{*}}|_{F^{*}}\leq\tau^{-1/2}|\theta-\hat{\theta}|_{H}+s\sqrt{d/n}\leq 2s\sqrt{d/n}$
We conclude that ${\mathcal{U}}(\sqrt{\tau}s)\subseteq{\mathcal{U}}^{*}(2s)$. Next, note that
$\begin{split}\delta_{3}(\sqrt{\tau}s)&=\sup_{\theta\in{\mathcal{U}}(\sqrt{\tau}s)}\|\nabla^{3}\ell(\theta)\|_{H}\leq\sup_{\theta\in{\mathcal{U}}^{*}(2s)}\|\nabla^{3}\ell(\theta)\|_{H}\\ &\leq\tau^{-3/2}\sup_{\theta\in{\mathcal{U}}^{*}(2s)}\|\nabla^{3}\ell(\theta)\|_{F^{*}}=\tau^{-3/2}\delta^{*}_{3}(2s).\end{split}$
Finally,
$\sqrt{\tau}s\delta_{3}(\sqrt{\tau}s)\sqrt{d/n}\leq(\sqrt{\tau}s)(\tau^{-3/2}\delta^{*}_{3}(2s))\sqrt{d/n}=\tau^{-1}s\delta^{*}_{3}(2s)\sqrt{d/n}\leq\tau^{-1}\lambda/4=1/2,$
proving (B.2). ∎
Appendix C Proofs from Section 4
Proof of Lemma 4.1.
Let $\bar{W}={F^{*}}^{-1/2}(\bar{Y}-\mathbb{E}\,_{\theta^{*}}[Y])$ and $\mathcal{N}$ be a $1/2$-net of the sphere $S^{d-1}$. Then
$|\bar{W}|\leq 2\sup_{u\in\mathcal{N}}u^{\intercal}\bar{W},$
and hence
$\begin{split}\mathbb{P}\big{(}|\bar{W}|&\geq 2s\sqrt{d/n}\big{)}\leq\mathbb{P}\left(\sup_{u\in\mathcal{N}}u^{\intercal}\bar{W}\geq s\sqrt{d/n}\right)\leq 5^{d}\sup_{|u|=1}\mathbb{P}\left(u^{\intercal}\bar{W}\geq s\sqrt{d/n}\right)\\ &=5^{d}\sup_{|u|_{F^{*}}=1}\mathbb{P}\left(u^{\intercal}(\bar{Y}-\mathbb{E}\,_{{\theta^{*}}}[Y])\geq s\sqrt{d/n}\right)\\ &=5^{d}\sup_{|u|_{F^{*}}=1}\mathbb{P}\left(\sum_{i=1}^{n}u^{\intercal}(Y_{i}-\mathbb{E}\,_{{\theta^{*}}}[Y_{i}])\geq s\sqrt{nd}\right)\\ &\leq 5^{d}e^{-\lambda sd}\sup_{|u|_{F^{*}}=1}\mathbb{E}\,_{{\theta^{*}}}\left[\,\mathrm{exp}\left(\lambda\sqrt{d/n}u^{\intercal}\left(Y_{1}-\mathbb{E}\,_{{\theta^{*}}}[Y_{1}]\right)\right)\right]^{n}.\end{split}$ (C.1)
In the second line, we used a union bound and the fact that $|\mathcal{N}|\leq(1+2/(1/2))^{d}=5^{d}$. In the last line we used Markov’s inequality. The exponential expectation is finite for all $\lambda$ such that ${\theta^{*}}+\lambda\sqrt{d/n}u\in\Theta$, as can be seen from the formula (C.2) below. Now, let $\tau=\lambda\sqrt{d/n}$. Using properties of exponential families, note that for some $|\xi|\leq\tau$, we have
$\begin{split}\mathbb{E}\,_{{\theta^{*}}}&\left[\,\mathrm{exp}\left(\tau u^{\intercal}\left(Y-\mathbb{E}\,_{{\theta^{*}}}[Y]\right)\right)\right]\\ &=\,\mathrm{exp}\left(\psi({\theta^{*}}+\tau u)-\psi({\theta^{*}})-\tau u^{\intercal}\psi({\theta^{*}})\right)=\,\mathrm{exp}\left(\frac{\tau^{2}}{2}u^{\intercal}\nabla^{2}\psi({\theta^{*}}+\xi u)u\right)\\ &\leq\,\mathrm{exp}\left(\frac{\tau^{2}}{2}\left[1+\lambda\delta^{*}_{3}(\lambda)\sqrt{d/n}\right]\right)=\,\mathrm{exp}\left(\frac{\lambda^{2}d}{2n}\left[1+\lambda\delta^{*}_{3}(\lambda)\sqrt{d/n}\right]\right)\end{split}$ (C.2)
We used that $\delta^{*}_{3}(\lambda)$ is the supremum of $\|\nabla^{3}\psi(\theta)\|_{F^{*}}$ over all $\theta$ such that $|\theta-{\theta^{*}}|_{F^{*}}\leq\lambda\sqrt{d/n}$, that $|u|_{F^{*}}=1$, and that $|\xi|\leq\lambda\sqrt{d/n}$. Substituting (C.2) into the last line of (C.1) gives
$\mathbb{P}(|\bar{W}|\geq s\sqrt{d/n})\leq\,\mathrm{exp}\left(d\left[\log 5-\lambda s+\frac{\lambda^{2}}{2}\left(1+\lambda\delta^{*}_{3}(\lambda)\sqrt{d/n}\right)\right]\right)$ (C.3)
Take $\lambda=s/2$, so that ${\theta^{*}}+(s/2)\sqrt{d/n}u\in{\mathcal{U}}^{*}(s/2)\subset\Theta$ for all $|u|_{F^{*}}=1$ by the assumption in the statement of the lemma. Therefore,  (C.2) is valid for this choice of $\lambda$. Using $(s/2)\delta^{*}_{3}(s/2)\sqrt{d/n}\leq 1$, also by assumption, the bound (C.3) leads to
$\mathbb{P}(|\bar{W}|\geq s\sqrt{d/n})\leq\,\mathrm{exp}\left(d\left[\log 5-\frac{s^{2}}{2}+\frac{s^{2}}{4}\right]\right)=\,\mathrm{exp}(d[\log 5-s^{2}/4]).$
Appendix D Proofs from Section 5
Proof of Lemma 5.5.
To prove the $\chi^{2}$ characterization of ${\mathcal{U}}^{*}(r)$, fix $\theta\in\mathbb{R}^{d}$ and let $\theta_{0}:=1-\theta^{\intercal}\mathds{1}$. Now note that
$\begin{split}|\theta-{\theta^{*}}|_{F^{*}}^{2}&=(\theta-{\theta^{*}})^{\intercal}{F^{*}}(\theta-{\theta^{*}})=\sum_{j=1}^{d}\frac{(\theta_{j}-\theta_{j}^{*})^{2}}{\theta_{j}^{*}}+\frac{((\theta-{\theta^{*}})^{\intercal}\mathds{1})^{2}}{\theta_{0}^{*}}\\ &=\sum_{j=1}^{d}\frac{(\theta_{j}-\theta_{j}^{*})^{2}}{\theta_{j}^{*}}+\frac{(\theta_{0}-\theta_{0}^{*})^{2}}{\theta_{0}^{*}}=\chi^{2}(\theta||{\theta^{*}}).\end{split}$ (D.1)
Hence ${\mathcal{U}}(r)=\{\theta\in\mathbb{R}^{d}\;:\;\chi^{2}(\theta||{\theta^{*}})\leq r^{2}d/n\}$. Note that in the first line above, ${\theta^{*}}$ denotes $(\theta_{1}^{*},\dots,\theta_{d}^{*})$. Next, suppose ${\mathcal{U}}^{*}(2r)\not\subseteq\Theta$. We will show that ${\theta_{\mathrm{min}}^{*}}\leq 4r^{2}d/n$. If ${\mathcal{U}}^{*}(2r)\not\subseteq\Theta$, then there exists $\theta\in{\mathcal{U}}^{*}(2r)$ and $j\in\{1,\dots,d\}$ such that $\theta_{j}\leq 0$ or $\theta_{j}\geq 1$. Suppose first that $\theta_{j}\leq 0$. Then
$\theta_{j}^{*}\leq\frac{(\theta_{j}-\theta_{j}^{*})^{2}}{\theta_{j}^{*}}\leq\chi^{2}(\theta||{\theta^{*}})\leq 4r^{2}d/n$ (D.2)
and hence ${\theta_{\mathrm{min}}^{*}}\leq 4r^{2}d/n$. Next, suppose $\theta_{j}\geq 1$. Then $\theta_{0}\leq 0$, and we can apply the above argument with $j=0$. Finally, fix $\theta\in{\mathcal{U}}^{*}(r)\subset\Theta$. Then the bound (5.6) gives
$\max_{j=0,\dots,d}\left|\frac{\theta_{j}}{\theta_{j}^{*}}-1\right|^{2}\leq\frac{\chi^{2}(\theta||{\theta^{*}})}{{\theta_{\mathrm{min}}^{*}}}\leq\frac{r^{2}d/n}{{\theta_{\mathrm{min}}^{*}}}\leq\frac{1}{4}$
by assumption on $r$, and hence $\theta_{j}\geq\theta_{j}^{*}/2\;\forall j=0,1,\dots,d$, as desired. ∎
Throughout the rest of the section, we use the shorthand notation
$\chi^{2}=\chi^{2}(\bar{N}||{\theta^{*}}).$
Proof of Lemma 5.7.
Using (4.3) of [6], we have
$\chi\leq C\sqrt{d/n}+C\sqrt{x/n}+C\frac{x}{n\sqrt{{\theta_{\mathrm{min}}^{*}}}}$ (D.3)
with probability at least $1-e^{-x}$ for any $x>0$. Now, let $x=dt^{2}$ for some $t\geq 1$ for which $\frac{t^{2}d}{n{\theta_{\mathrm{min}}^{*}}}\leq 1$. Since $t\geq 1$ we can upper bound (D.3) as
$\chi\leq C\sqrt{t^{2}d/n}+C\frac{t^{2}d}{n\sqrt{\theta_{\mathrm{min}}^{*}}}=C\sqrt{\frac{t^{2}d}{n}}\left(1+\sqrt{\frac{t^{2}d}{n{\theta_{\mathrm{min}}^{*}}}}\right)\leq C_{1}\sqrt{\frac{t^{2}d}{n}},$ (D.4)
using the assumption $\frac{t^{2}d}{n{\theta_{\mathrm{min}}^{*}}}\leq 1$. The bound (D.4) holds with probability at least $1-e^{-t^{2}d}$. We now choose $t=s/C_{1}$, which satisfies $t\geq 1$ and $\frac{t^{2}d}{n{\theta_{\mathrm{min}}^{*}}}\leq 1$ if $s\geq C_{1}$ and $s^{2}d/n\leq C_{1}^{2}{\theta_{\mathrm{min}}^{*}}$. Thus $E(s)=\{\chi\leq s\sqrt{d/n}\}$ holds with probability at least $1-e^{-t^{2}d}=1-e^{-s^{2}d/C_{1}^{2}}$. ∎
Remark D.1.
Since we have assumed $s^{2}d/n\leq C_{1}^{2}{\theta_{\mathrm{min}}^{*}}$, note that we also have $\chi^{2}/{\theta_{\mathrm{min}}^{*}}\leq s^{2}d/(n{\theta_{\mathrm{min}}^{*}})\leq C_{1}^{2}$ on $E(s)$.
Proof of Lemma 5.8.
First recall that since $(2s)^{2}d/n\leq{\theta_{\mathrm{min}}^{*}}/4$, Lemma 5.5 shows that
$\frac{\theta_{j}^{*}}{\theta_{j}}\leq 2\quad\forall j=0,\dots,d\quad\forall\theta\in{\mathcal{U}}^{*}(2s).$ (D.5)
Now, fix $\theta\in{\mathcal{U}}^{*}(2s)$. We upper bound $\|\nabla^{3}\ell(\theta)\|_{F^{*}}$, which is given by the maximum of
$\begin{split}\left\langle\nabla^{3}\ell(\theta),u^{\otimes 3}\right\rangle=-2\sum_{j=1}^{d}\frac{\bar{N}_{j}}{\theta_{j}^{3}}u_{j}^{3}+2\frac{\bar{N}_{0}}{\theta_{0}^{3}}(u^{\intercal}\mathds{1})^{3}=-2\sum_{j=0}^{d}\frac{\bar{N}_{j}}{\theta_{j}^{3}}u_{j}^{3}\end{split}$ (D.6)
over all $u$ such that
$1=u^{\intercal}{F^{*}}u=\sum_{j=1}^{d}\frac{u_{j}^{2}}{\theta_{j}^{*}}+\frac{(u^{\intercal}\mathds{1})^{2}}{\theta_{0}^{*}}=\sum_{j=0}^{d}\frac{u_{j}^{2}}{\theta_{j}^{*}},$ (D.7)
where $u_{0}:=-\sum_{j=1}^{d}u_{j}$. Dropping this constraint on the sum of the $u_{j}$ and using Lagrange multipliers, we see that the maximum is achieved at a $u$ such that $u_{j}=0$ for $j\in I$ and $u_{j}=\lambda\theta_{j}^{3}/(\theta_{j}^{*}\bar{N}_{j})$ for $j\in\{0,\dots,d\}\setminus I$, where $I$ is some non-empty index set. The constraint (D.7) reduces to $\lambda^{2}S=1$ and the objective (D.6) reduces to $-2\lambda^{3}S$, where $S=\sum_{i\in I}\frac{\theta_{j}^{6}}{(\theta_{j}^{*})^{3}\bar{N}_{j}^{2}}$. Thus $\lambda=-1/\sqrt{S}$, and the objective (D.6) is given by $-2\lambda^{3}S=2/\sqrt{S}$, which is maximized when $S=\min_{j=0,\dots,d}\frac{\theta_{j}^{6}}{(\theta_{j}^{*})^{3}\bar{N}_{j}^{2}}$. Thus
$\|\nabla^{3}\ell(\theta)\|_{F^{*}}\leq 2\max_{j=0,\dots,d}\frac{\bar{N}_{j}(\theta_{j}^{*})^{3/2}}{\theta_{j}^{3}}$ (D.8)
We now take the supremum over all $\theta\in{\mathcal{U}}^{*}(2s)$ and apply (D.5) to get that
$\delta^{*}_{3}(2s)=\sup_{\theta\in{\mathcal{U}}^{*}(2s)}\|\nabla^{3}\ell(\theta)\|_{F^{*}}\leq C\max_{j=0,\dots,d}\frac{1}{\sqrt{\theta_{j}^{*}}}\frac{\bar{N}_{j}}{\theta_{j}^{*}}\leq\frac{C}{\sqrt{\theta_{\mathrm{min}}^{*}}}\left(1+\frac{\chi}{\sqrt{\theta_{\mathrm{min}}^{*}}}\right).$ (D.9)
The last inequality follows by (5.6) applied with $\theta=\bar{N}$. Finally, as noted in Remark D.1, we have that $\chi/\sqrt{{\theta_{\mathrm{min}}^{*}}}$ is bounded by an absolute constant on $E(s)$. This concludes the proof. ∎
Calculation from Remark 5.9.
In the case that $\bar{N}_{j}=\theta_{j}^{*}$, we have as in the above proof that
$\|\nabla^{3}\ell({\theta^{*}})\|_{F^{*}}=\sup\left\{-2\sum_{j=0}^{d}\frac{u_{j}^{3}}{{\theta_{j}^{*}}^{2}}\;:\;\sum_{j=0}^{d}u_{j}=0,\sum_{j=0}^{d}\frac{u_{j}^{2}}{\theta_{j}^{*}}=1\right\}$ (D.10)
Using Lagrange multipliers we get that $u_{j}=c_{1}\theta_{j}^{*}$ for all $j$ in some nonempty index set $I$, and $u_{j}=c_{2}\theta_{j}^{*}$ for all other $j$. If $P=\sum_{j\in I}\theta_{j}^{*}$, the constraints reduce to $c_{1}P+c_{2}(1-P)=0$ and $c_{1}^{2}P+c_{2}^{2}P=1$. Solving for $c_{1},c_{2}$ we get $c_{1}=-\sqrt{P/(1-P)}$ and $c_{2}=\sqrt{(1-P)/P}$. The objective value is then equal to $2[(1-P)^{3/2}/\sqrt{P}-P^{3/2}/\sqrt{1-P}]=2(1-2P)/\sqrt{P(1-P)}$. This is maximized when $P$ is smallest, which is $P={\theta_{\mathrm{min}}^{*}}$. This yields
$\|\nabla^{3}\ell({\theta^{*}})\|_{F^{*}}=2\frac{1-2{\theta_{\mathrm{min}}^{*}}}{\sqrt{{\theta_{\mathrm{min}}^{*}}}\sqrt{1-{\theta_{\mathrm{min}}^{*}}}}$ (D.11)
Proof of Lemma 5.10.
We have
$\begin{split}\|\nabla^{2}\ell({\theta^{*}})-{F^{*}}\|_{F^{*}}&=\sup_{u\neq 0}\frac{\sum_{j=1}^{d}\frac{\bar{N}_{j}-\theta_{j}^{*}}{(\theta_{j}^{*})^{2}}u_{j}^{2}+\frac{\bar{N}_{0}-\theta_{0}^{*}}{(\theta_{0}^{*})^{2}}(\mathds{1}^{\intercal}u)^{2}}{\sum_{j=1}^{d}\frac{1}{\theta_{j}^{*}}u_{j}^{2}+\frac{1}{\theta_{0}^{*}}(\mathds{1}^{\intercal}u)^{2}}\\ &\leq\max_{j=0,1,\dots,d}\frac{|\bar{N}_{j}-\theta_{j}^{*}|}{\theta_{j}^{*}}\leq\frac{\chi}{\sqrt{\theta_{\mathrm{min}}^{*}}}\leq\sqrt{\frac{s^{2}d}{n{\theta_{\mathrm{min}}^{*}}}},\end{split}$ (D.12)
using (5.6) to get the second-to-last inequality, and that $\chi\leq s\sqrt{d/n}$ on $E(s)$ to get the last inequality. ∎
Appendix E Proofs from Section 6 (Logistic Regression)
Proof of Lemma 6.2.
Let $r=2\lambda s$ and $Q=\frac{1}{n}\sum_{i=1}^{n}(Y_{i}-\mathbb{E}\,[Y_{i}\mid X_{i}])X_{i}$. We are interested in bounding from above the tail probability $\mathbb{P}(\|Q\|_{S}\geq r\sqrt{d/n})$. Let $\mathcal{N}$ be a $1/2$-net of the sphere in $\mathbb{R}^{d}$, i.e. a collection of unit norm vectors $u$ such that $\inf_{u\in\mathcal{N}}|u-w|\leq 1/2$ for all $|w|=1$. By standard arguments, we can take $\mathcal{N}$ to have at most $5^{d}$ elements. Next, let $\mathcal{N}_{S}:=\{S^{-1/2}u\mid u\in\mathcal{N}\}$. Then
$\|Q\|_{S}\leq 2\sup_{u\in\mathcal{N}_{S}}u^{\intercal}Q,$
and hence
$\mathbb{P}(\|Q\|_{S}\geq r\sqrt{d/n})\leq\mathbb{P}(\sup_{u\in\mathcal{N}_{S}}u^{\intercal}Q\geq(r/2)\sqrt{d/n})\leq 5^{d}\sup_{|u|_{S}=1}\mathbb{P}(u^{\intercal}Q\geq(r/2)\sqrt{d/n}),$ (E.1)
where we have used a union bound and the fact that $|\mathcal{N}_{S}|\leq 5^{d}$. Now, fix $|u|_{S}=1$. Then
$\begin{split}\mathbb{P}(u^{\intercal}Q\geq(r/2)\sqrt{d/n})&=\mathbb{E}\,\left[\mathbb{P}\left(u^{\intercal}Q\geq(r/2)\sqrt{d/n}\mid\{X_{i}\}_{i=1}^{n}\right)\right]\\ &=\mathbb{E}\,\left[\mathbb{P}\left(\sum_{i=1}^{n}(Y_{i}-\mathbb{E}\,[Y_{i}\mid X_{i}])u^{\intercal}X_{i}\geq(r/2)\sqrt{nd}\mid\{X_{i}\}_{i=1}^{n}\right)\right]\\ &\leq\mathbb{E}\,\left[\,\mathrm{exp}\left(-\frac{2(r/2)^{2}nd}{\sum_{i=1}^{n}(u^{\intercal}X_{i})^{2}}\right)\right]=\mathbb{E}\,\left[\,\mathrm{exp}\left(-\frac{r^{2}nd}{2|Z|^{2}}\right)\right],\end{split}$ (E.2)
where $Z\sim\mathcal{N}(0,I_{n})$. We have used Hoeffding’s inequality to get the bound in the third line. The final equality holds by the fact that $(u^{\intercal}X_{i})_{i=1}^{n}$ is a standard normal on $\mathbb{R}^{n}$. Indeed, the $X_{i}$ are independent and we can write $u=S^{-1/2}w$ for some unit vector $w$, so that $u^{\intercal}X_{i}=w^{\intercal}S^{-1/2}X_{i}\sim\mathcal{N}(0,1)$. Now, we have
$\mathbb{E}\,\left[\,\mathrm{exp}\left(-\frac{r^{2}nd}{2|Z|^{2}}\right)\right]\leq\mathbb{P}(|Z|\geq 2\sqrt{n})+\,\mathrm{exp}\left(-\frac{r^{2}nd}{2(2\sqrt{n})^{2}}\right)\leq e^{-n/2}+e^{-r^{2}d/8}$ (E.3)
Combining (E.1), (E.2), and (E.3), and using $5\leq e^{2}$, we conclude that
$\mathbb{P}(\|Q\|_{S}\geq r\sqrt{d/n})\leq e^{2d-n/2}+e^{d(2-r^{2}/8)}\leq e^{-n/4}+e^{-d(\lambda s)^{2}/4}.$
The second inequality uses that $r=2s\lambda$, $(\lambda s)^{2}\geq 8$, and $d/n\leq 1/8$. ∎
Proof of (6.5) via Theorem 6.3.
First note that $\sup_{|u|_{S}=1}\frac{1}{n}\sum_{i=1}^{n}|u^{\intercal}X_{i}|^{3}\stackrel{{\scriptstyle d}}{{=}}\sup_{|u|=1}\frac{1}{n}\sum_{i=1}^{n}|u^{\intercal}Z_{i}|^{3},$ so it suffices to show the latter quantity is bounded with high probability. We apply Theorem 6.3 with $p=3$. Let $N=n/d^{1.5}$ and $t=bN^{1/8}$, for $b$ to be determined. Then
$\begin{split}t^{6}[\log(2t^{2})]^{4}=b^{6}N^{3/4}[\log(2b^{2}N^{1/4})]^{4}\leq Cb^{6}N^{3/4}[\log^{4}(2b^{2})+\log^{4}(N)]\leq CN^{3/4}(b^{2}+b^{6}\log^{4}(N)).\end{split}$
The last inequality is true for $b$ small enough. Therefore, $Ct^{6}[\log(2t^{2})]^{4}\leq N$ is satisfied if $C(b^{2}+b^{6}\log^{4}N)\leq N^{1/4}$. This holds if $N\geq 1$ and $b$ is small enough. Note that we can write $t$ as $t=bN^{1/8}=bn^{1/8}d^{-3/16}$, so that $Ct\sqrt{d}=Cn^{1/8}d^{5/16}\geq C(nd)^{1/8}$, where $b$ got absorbed into $C$.
Thus we conclude that if $n\geq d^{3/2}$ then $\sup_{|u|=1}\frac{1}{n}\sum_{i=1}^{n}|u^{\intercal}Z_{i}|^{3}\leq C$ with probability at least $1-e^{-C(nd)^{1/8}}$ as desired. ∎
Appendix F Auxiliary
Lemma F.1.
Let ${F^{*}},H$ be two symmetric positive definite matrices in $\mathbb{R}^{d\times d}$ such that $H\succeq\lambda{F^{*}}$. Also, let $u\in\mathbb{R}^{d}$ and $T$ be a symmetric $k$ tensor for $k\geq 1$. Then
$|u|_{H}\geq\lambda^{1/2}|u|_{F^{*}},\quad\|T\|_{H}\leq\lambda^{-k/2}\|T\|_{F^{*}}$ (F.1)
Proof.
The first bound is by definition. For the second bound, we use the first bound to get
$\|T\|_{H}=\sup_{u\neq 0}\frac{\langle T,u^{\otimes k}\rangle}{(u^{\intercal}Hu)^{k/2}}\leq\lambda^{-k/2}\sup_{u\neq 0}\frac{\langle T,u^{\otimes k}\rangle}{(u^{\intercal}{F^{*}}u)^{k/2}}=\lambda^{-k/2}\|T\|_{F^{*}}$ (F.2)
Lemma F.2.
Suppose $\tau_{1}{F^{*}}\preceq A_{1}$, $\tau_{2}{F^{*}}\preceq A_{2}$, and $\|A_{1}-A_{2}\|_{F^{*}}\leq\epsilon$. Then
$\mathrm{TV}\left(\mathcal{N}(x,(nA_{1})^{-1}),\;\mathcal{N}(x,(nA_{2})^{-1})\right)\leq\frac{\epsilon}{2}(\tau_{1}^{-1}+\tau_{2}^{-1})\sqrt{d}.$ (F.3)
Proof.
First, we use Pinsker’s inequality and the formula for the KL divergence between two Gaussians:
$\begin{split}\mathrm{TV}&\left(\mathcal{N}(x,(nA_{1})^{-1}),\;\mathcal{N}(x,(nA_{2})^{-1})\right)^{2}\leq\frac{1}{2}\mathrm{KL}\left(\,\mathcal{N}(x,(nA_{1})^{-1})\;||\;\mathcal{N}(x,(nA_{2})^{-1})\right)\\ &=\frac{1}{2}\mathrm{Tr}(A_{2}A_{1}^{-1}-I_{d})-\frac{1}{2}\log\det(A_{2}A_{1}^{-1})=\frac{1}{2}\mathrm{Tr}(B_{2}B_{1}^{-1}-I_{d})-\frac{1}{2}\log\det(B_{2}B_{1}^{-1})\end{split}$ (F.4)
where $B_{i}={F^{*}}^{-1/2}A_{i}{F^{*}}^{-1/2}$, $i=1,2$. Note that
$\|B_{1}-B_{2}\|=\|A_{1}-A_{2}\|_{F^{*}}\leq\epsilon,\qquad\lambda_{\min}(B_{i})\geq\tau_{i}.$ (F.5)
Now, for $\lambda>0$, a Taylor expansion of the logarithm gives
$\begin{split}\lambda-1-\log\lambda=\frac{(\lambda-1)^{2}}{2\xi^{2}}\leq\frac{(\lambda-1)^{2}}{2}+\frac{(\lambda-1)^{2}}{2\lambda^{2}}=\frac{(\lambda-1)^{2}}{2}+\frac{(\lambda^{-1}-1)^{2}}{2}\end{split}$
In the equality, $\xi$ is a point between $1$ and $\lambda$. It follows that for a positive definite symmetric matrix $M\in\mathbb{R}^{d\times d}$, we have
$\begin{split}\mathrm{Tr}(M-I_{d})-\log\det M&\leq\frac{d}{2}\left(\|M-I_{d}\|^{2}+\|M^{-1}-I_{d}\|^{2}\right).\end{split}$ (F.6)
Using (F.6) in (F.4), we get
$\begin{split}\mathrm{TV}\big{(}&\mathcal{N}(x,(nA_{1})^{-1}),\;\mathcal{N}(x,(nA_{2})^{-1})\big{)}^{2}\leq\frac{d}{4}(\|B_{2}B_{1}^{-1}-I_{d}\|^{2}+\|B_{1}B_{2}^{-1}-I_{d}\|^{2})\\ &\leq\frac{d}{4}\|B_{2}-B_{1}\|^{2}(\lambda_{\min}(B_{1})^{-2}+\lambda_{\min}(B_{2})^{-2})\leq\frac{d}{4}\epsilon^{2}(\tau_{1}^{-2}+\tau_{2}^{-2}).\end{split}$
The last line uses (F.5). Taking the square root concludes the proof. ∎
Lemma F.3 (Lemma E.3 of [14] with $p=0$).
If $ab>1$, then
[…]
Lemma F.4 (Modified from Lemma 1.3 in Chapter XIV of [15]).
Let $U\subset\mathbb{R}^{d}$ be open, and let $f:U\to\mathbb{R}^{d}$ be of class $C^{1}$. Assume that $f(0)=0$ and $\nabla f(0)=I_{d}$. Let $r>0$ be such that $\bar{B}_{r}(0)\subset U$. If
$|\nabla f(x)-\nabla f(0)|\leq\frac{1}{4},\qquad\forall|x|\leq r,$
then $f$ maps $\bar{B}_{r}(0)$ bijectively onto $\bar{B}_{2r}(0)$.
Now let $h:\mathbb{R}^{d}\to\mathbb{R}$ be of class $C^{2}$, such that $H=\nabla^{2}h(y_{0})$ is symmetric positive definite. Define
$\tilde{h}(x)=H^{-1/2}\left(\nabla h\left(y_{0}+H^{-1/2}x\right)-\nabla h(y_{0})\right).$
Then $\tilde{h}(0)=0$ and $\nabla\tilde{h}(0)=I_{d}$, and one can check that $|\nabla\tilde{h}(x)-\nabla\tilde{h}(0)|\leq 1/4$ for all $x\in\bar{B}_{r}(0)$ if and only if
$\|\nabla^{2}h(y)-\nabla^{2}h(y_{0})\|_{H}\leq 1/4,\quad\forall|y-y_{0}|_{H}\leq r.$ (F.8)
We conclude that if (F.8) is satisfied, then $\nabla\tilde{h}$ maps $\bar{B}_{r}(0)$ bijectively onto $\bar{B}_{2r}(0)$ and hence $\nabla h$ maps $y_{0}+H^{-1/2}\bar{B}_{r}(0)$ bijectively onto $\nabla h(y_{0})+H^{1/2}\bar{B}_{2r}(0)$. We have proved the following corollary.
Corollary F.5.
Let $h:\mathbb{R}^{d}\to\mathbb{R}$ be $C^{2}$ such that $H=\nabla^{2}h(y_{0})$ is symmetric positive definite. Suppose (F.8) holds, and suppose $z$ is such that
$\|z-\nabla h(y_{0})\|_{H}=|H^{-1/2}(z-\nabla h(y_{0}))|\leq 2r.$
Then there exists $y$ such that $|y-y_{0}|_{H}\leq r$ and $\nabla h(y)=z$.
Next, it will be convenient to use a weighting matrix ${F^{*}}$ that does not exactly equal $H=\nabla^{2}h(y_{0})$. We will also only need the case $z=0$.
Corollary F.6.
Let $h:\mathbb{R}^{d}\to\mathbb{R}$ be $C^{2}$ such that $H=\nabla^{2}h(y_{0})$ is symmetric positive definite, and let ${F^{*}}$ be a symmetric positive definite matrix such that $H\succeq\lambda{F^{*}}$. Suppose
$\begin{split}\|\nabla^{2}h(y)-\nabla^{2}h(y_{0})\|_{{F^{*}}}&\leq\frac{\lambda}{4}\quad\forall|y-y_{0}|_{{F^{*}}}\leq q,\\ |{F^{*}}^{-1/2}\nabla h(y_{0})|&\leq 2\lambda q\end{split}$
for some $q\geq 0$. Then there exists $y$ such that $|y-y_{0}|_{F^{*}}\leq q$ and $\nabla h(y)=0$.
Proof.
If $|y-y_{0}|_{H}\leq r=\sqrt{\lambda}q$, then $|y-y_{0}|_{F^{*}}\leq q$ and hence $\|\nabla^{2}h(y)-\nabla^{2}h(y_{0})\|_{{F^{*}}}\leq\lambda/4$, which implies $\|\nabla^{2}h(y)-\nabla^{2}h(y_{0})\|_{H}\leq 1/4$. Also, note that $|H^{-1/2}\nabla h(y_{0})|\leq 2\sqrt{\lambda}q=2r$. Thus the assumptions of Corollary F.5 are satisfied, so that there exists $y$ such that $|y-y_{0}|_{H}\leq r$ and $\nabla h(y)=0$. But then $|y-y_{0}|_{F^{*}}\leq r/\sqrt{\lambda}=q$, as desired. ∎